### v1

# A Statistical Machine Learning Approach for Adapting Reduced-Order Models using Projected Gaussian Process

Abstract
Projection-based model reduction is among the most widely adopted approaches for constructing parametric Reduced-Order Models (ROMs). Utilizing snapshot data from solving full-order governing equations, the Proper Orthogonal Decomposition (POD) computes the optimal basis modes that span a low-dimensional subspace where the ROM resides. Because a governing equation is often parameterized by a set of parameters, challenges immediately arise when one would like to investigate how systems behave differently over the parameter space (in design, diagnosis, control, uncertainty quantification and real-time operations). In this case, the optimal basis needs to be efficiently updated so as to adapt ROM that can accurately capture the variation of a system’s behavior over its parameter space. In this paper, we propose a Projected Gaussian Process (pGP) model and formulate the problem of adapting POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping (injective) from the parameter space to the Grassmann Manifold that contains the optimal vector subspaces. To establish such a relationship, a mapping is found between the Euclidean space and the horizontal space of an orthogonal matrix that spans a reference subspace in the Grassmann Manifold. Then, a second mapping from the horizontal space to the Grassmann Manifold is established through the Exponential/Logarithm maps between the manifold and its tangent space. Finally, given a new parameter, the conditional distribution of a vector can be found in the Euclidean space using the Gaussian Process (GP) regression, and such a distribution is then projected to the Grassmann Manifold that enables us to find the optimal subspace, i.e., POD basis, for the new parameter. Compared with existing interpolation method, the proposed statistical learning approach allows us to optimally estimate (or tune) model parameters given data (i.e., the prediction/interpolation becomes problem-specific), and quantify the uncertainty associated with the prediction. Numerical examples are presented to demonstrate the advantages of the proposed pGP for adapting POD basis against parameter changes.
keywords:
Reduced-Order Models, Proper Orthogonal Decomposition, Gaussian Process, Projected Gaussian Process, Grassmann Manifold
1 Introduction
1.1 Background and Problem Statement
Consider the diffusion of chemical species, aeroelasticity of aircraft wings, flow of vehicles and predator-prey interactions, dynamics of such complex systems is typically described by a set of Partial Differential Equations (PDEs). Very often, the dimensions of these problems are extremely high due to the spatial discretization of the mathematical models (e.g., Computational Fluid Dynamics models or Finite Element Analysis with very large numbers of degrees of freedom). The prohibitive computational cost prevents us from repeatedly running numerical solvers under different parameter settings; often needed for design, diagnosis, prediction, uncertainty quantification and control problems[24, 25, 29]. Hence, when there exist low-dimensional patterns embedded in high-dimensional systems[16], Reduced-Order Models (ROMs) become indispensable in capturing the dominate behaviors of complex systems. Compared with data-driven surrogate models or meta models[3], ROM often preserve important system physics, such as stability[26, 4], passivity[31, 9], polynomial nonlinearities[27], etc.
Projection-based model reduction is among the most widely adopted approaches for constructing ROMs[16, 27]. The central idea behind projection-based ROM is to project a high-dimensional governing equation to an optimal low-dimensional vector space spanned by orthogonal bases. Although searching for such bases may seem to be nontrivial considering the computational speed, accuracy and complex boundary conditions, the Proper Orthogonal Decomposition (POD)—applied to curves in Hilbert spaces $\mathcal{H}$ of infinite dimension—provides a computationally efficient method for approximating data generated from full-order systems in an $\ell_{2}$ sense[6]. This can be efficiently done by performing the thin Singular Value Decomposition (SVD) of the snapshot data matrix[16].
However, a governing equation is often parameterized by a set of parameters, and challenges immediately arise when one would like to investigate how systems behave differently over the parameter space (of the governing equation). In this case, the optimal basis needs to be updated so that the ROM can accurately capture the variation of a system’s behavior over its parameter space. As pointed out by [10, 32], the robustness and accuracy of an ROM are related to the reference parameters from which POD modes are derived. The accuracy of ROM determines the performance of any downstream machine learning models built on the low-dimensional ROM.
To elaborate, consider a governing equation with parameters $\bm{\theta}=(\theta_{1},\theta_{2},...,\theta_{d})\in\mathbb{P}\subset\mathbb{R}^{d}$, where $\mathbb{P}$ is the parameter space. Given $\bm{\theta}$, the solution of the system, $x(t,\bm{s};\bm{\theta})$, is a space-time field,
(1) $(t,\bm{s})\in[0,T]\times\mathbb{S}\mapsto x(t,\bm{s};\bm{\theta})$
where $\mathbb{S}$ is the spatial domain and $T>0$. Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),x(t,\bm{s}_{2};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}\in\mathbb{R}^{n}$ (say, using the Finite Element method), one can obtain the so-called snapshot data matrix, $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$, where $n_{T}$ is the number of temporal points. The POD method finds the optimal $r$-dimensional ($r\ll n$) vector subspace $\mathbb{W}\subset\mathbb{R}^{n}$ and an orthogonal projection, $\mathcal{W}:\mathbb{R}^{n}\rightarrow\mathbb{W}$ such that the following distance is minimized[12]
(2) $\text{min}_{\mathcal{W}}\sum_{i=1}^{n_{T}}||\bm{x}(t_{i};\bm{\theta})-\mathcal{W}(\bm{x}(t_{i};\bm{\theta}))||^{2}$
[…]
Hence, we see both the advantages and limitations of the POD approach above. The optimal POD basis, which is obtained from the snapshot data generated at the parameter setting $\bm{\theta}$, is optimal and only optimal for parameter $\bm{\theta}$. As shown in Figure 1, if data are already available from numerically solving the full-order governing equation repeatedly at a set of pre-determined parameter settings, it is possible to obtain the optimal basis for each parameter setting using the existing approach. However, for any new parameter in the parameter space, new basis modes (which is optimal or near optimal for the new parameter) is needed so that one can adapt the ROM according to parameter changes. The lack of robustness against parameter change limits the applicability of ROMs for design, diagnosis, control, uncertainty quantification and real-time operations (e.g., digital twins), where parameter changes are common[1, 24, 25, 29].
In this paper, we formulate the problem of adapting POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping between the parameter $\bm{\theta}$ and POD basis $\bm{\Phi}$ that spans a low-dimensional vector subspace. This subspace regression problem can be stated as follows:
Given $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$ for $\bm{\theta}_{i}\in\mathbb{P}$, the problem is to learn a mapping $\mathcal{T}:\mathbb{P}\rightarrow\mathcal{ST}(r,n)$ from the parameter space $\mathbb{P}\subset\mathbb{R}^{d}$ to a compact Stiefel manifold of matrices with orthogonal columns, which enables the prediction of the POD basis matrix $\bm{\Phi}^{*}$ given a new parameter $\bm{\theta}^{*}$ without numerically solving the computationally expensive full-order governing equation at $\bm{\theta}^{*}$ (see Figure 1).
We address this problem by proposing a supervised statistical learning approach, known as the Projected Gaussian Process (pGP) regression. As shown in Figure 1, when (training) data are available by numerically solving the full-order governing equation repeatedly at a (usually small) number of parameter settings $\{\bm{\theta}_{i}\}_{i=1}^{k}$, it is possible to obtain the optimal basis $\{\bm{\Phi}_{i}\}_{i=1}^{k}$ at each parameter setting, separately, using the existing approach. Once the proposed pGP regression learns the mapping between parameters and POD bases, a new basis can be generated given the new parameters, thus adapting the ROM for parameter changes (see Sec.3).
Figure 1: A supervised statistical learning problem (based on the proposed Projected Gaussian Process) that establishes the mapping between parameters and POD basis modes so that ROMs can be adapted for parameter changes.
1.2 Literature Review
Various strategies have been proposed to adapt ROM for parameter changes. Figure 2 presents one possible way to group existing approaches into four categories. This figure is used to facilitate the explanations of different approaches as well as how the proposed approach is built upon prior work. The boundaries that separate different approaches may change if we examine these methods from different perspectives.
$\bullet$ Strategy I involves pre-computing a database, i.e., a library, of reduced-order bases[1, 15]. For applications where this approach is feasible, ROM can be updated in real time over the set of parameters included in the database.
$\bullet$ Strategy II computes the first-order total derivatives of the POD modes with respective to a parameter[8], and a new POD basis, given a small change of the parameters from a base point, can be found through the first-order expansion in the parametric space.
$\bullet$ Strategy III involves different approaches that aim to interpolate or predict new POD bases from the observed parameter-POD basis pairs. In particular, the POD basis interpolation on Grassmann Manifolds have been extensively investigated[2, 30, 6, 23]. This approach is based on an important result that a subspace spanned by a POD basis is a point of the Grassmann Manifold that consists of $r$-dimensional subspaces in $\mathbb{R}^{n}$, where $n$ and $r$ are respectively the original and reduced-order dimensions. Hence, a tangent space, attached to a given base point in the Grassmann Manifold, can be constructed, and the subspaces from the Grassmann Manifold can be projected to the tangent space through a logarithmic map. After this critical step, interpolation can be done on the tangent space (such as the Lagrangian interpolation[2], inverse distance weighting[23], or Gaussian Process[13]), and the interpolated vector on the tangent space is eventually mapped back to the Grassmann Manifold through an exponential map. Hence, this method depends on the choice of the base point upon which the tangent space is attached to, and naturally requires the injective radius of the exponential map to be sufficiently large so that the approach is stable[12]. In contrast, the subspace angle interpolation approach directly computes the geometric distance between two subspaces by finding the principal angels between two POD bases[34]. The principal angels can be viewed as a series of angles required to rotate one subspace to form another, and such a distance metric is used to interpolate subspaces.
Figure 2: Existing strategies to adapt ROM for parameter changes
$\bullet$ Strategy IV involves statistical machine learning to learn the mapping between parameters and POD bases (or the subspaces spanned by POD bases). Leveraging the key ideas from both the POD basis interpolation on Grassmann Manifold[2] and subspace angle interpolation[34], a regression tree-based approach on Grassmann Manifold was proposed in [18]. Following this approach, a tree is grown by repeatedly splitting the tree node (i.e., a binary partition the parameter space) to maximize the Riemannian distance between two subspaces (based on the principals angles) spanned by the predicted POD bases on the left and right daughter nodes. As a result, this tree-based approach effectively divides the parameter space into sub-regions, and find the optimal POD basis for each sub-region (which could generate useful insights on how a system’s behavior varies over the parameter space). This approach successfully combines the POD basis prediction problem into the classical framework of regression trees, and appears to be more stable and conservative than the existing POD basis interpolation on Grassmann Manifolds (as the tree-based method only updates the POD basis when the parameter jumps from one sub-region to another in the parameter space). We also note that, this approach is useful for finding the POD basis when the dimension of the parameter space is high and the system behaviors have different sensitivity levels against different parameters. For those less important parameters, the tree-based approach effectively exclude those parameters from being used to split the tree (note that, the existing interpolation method does not automatically do this).
Other statistical learning approaches involve the use of Gaussian Process (GP). [13] directly used GP to capture the relationship between parameters and the matrices that determine the vectors on the tangent space of the Grassmann Manifold. Hence, given a new parameter setting, GP is used to predict a new vector in the tangent space, which yields the predicted subspace on the Grassmann Manifold through an exponential map. Note that, because GP performs prediction through a weighted linear combination of responses, this approach is similar to the existing Lagrangian POD basis interpolation on Grassmann Manifolds but essentially with different interpolation weights. For the reasons noted in Sec.3, we refer this approach as the native GP approach following the terminology in [21]. [20] proposed an approach known as Common POD (cPOD). In the context of engine injector design, a common POD basis is found for some base design geometry, the POD bases for other geometries are obtained by rescaling the common POD through a linear map. This approach requires sound domain knowledge of the system under consideration to determine the base design and establish the linear map between POD bases from the common POD. [35] proposed a GP subspace regression where the central idea is to directly construct a matrix-variate GP for the matrix bases such that the approach does not hinge on the differential geometric structures of the Gassmann manifold.
1.3 Contributions of this Paper
The prior work in the literature has shown the great potential of leveraging machine learning for adapting projection-based ROMs. The main contribution of this paper is to propose a new supervised statistical learning method, known as the Projected Gaussian Process (pGP) regression, that enables the learning of the mapping between parameters and POD bases. To our best knowledge, this method is not yet available in the literature. Computer code is also available as Supplementary Materials.
As a statistical machine learning approach, the proposed pGP provides some critical capabilities much needed for adapting ROMs against parameter changes:
i). Unlike existing interpolation method (e.g., Lagrangian interpolation), the statistical learning approach allows us to optimally estimate (or tune) model parameters based on the training data (i.e., the prediction becomes problem-specific). In contrast, when the Lagrangian interpolation is used, the Lagrange bases are computed from pre-existing equations which are independent of the problem or data of interest. Because the pGP model is always trained and optimized given the training dataset, such a flexibility potentially provides the opportunity to further improve the accuracy of ROM for the specific problem at hand (as shown in the numerical examples)
ii). The proposed pGP, as a statistical learning approach, provides the capability of parameter selection when constructing the mapping between parameters and POD bases. This capability is largely missing in the literature but is particularly important for systems with high-dimensional parameter space, i.e., not all parameters are equally important in terms of how they change the POD basis modes. If relevant parameters can be identified, one only needs to explore a reduced parameter space, with a lower dimension, when investigating how system dynamics varies in the parameter space.
iii). Due to the parameter estimation/selection capabilities, the pGP approach naturally involves the use of global POD basis as its special case. When no parameter is identified as significant, the predicted POD basis matrices for different parameters, under the proposed pGP framework, become i.i.d. samples from a matrix-variate Gaussian distribution with the global POD basis matrix being the mean of this distribution (as shown in numerical Example II).
iv). The proposed pGP not only provides the point prediction of the POD basis, but also scientifically quantifies the uncertainty associated with the predicted POD basis as well as the output generated by ROMs. Such a capability of uncertainty quantification facilitates decision-making under uncertainty, and provides a critical link (as a future research path) from the proposed approach to a large body of literature on statistical experimental design and deep reinforcement learning for more efficient construction of ROMs by sequentially exploring the solution space of full-order models[14, 7].
The paper is organized as follows. In Section 2, we provide the preliminaries on POD basis interpolation. The proposed pGP approach is presented in Section 3. Comprehensive numerical examples and discussions are presented in Section 4. Section 5 concludes the paper.
2 Preliminaries on POD basis interpolation
This section provides preliminaries on the interpolation problem of POD basis modes. As discussed in Section 1.1, due to the computational cost associated with solving full-order governing equations, it is only possible to obtain the snapshot matrices for a relatively small number of parameters, $\bm{\Lambda}=\{\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{k}\}$, and obtain the optimal POD basis $\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}$ for each parameter in $\bm{\Lambda}$. Hence, for any new parameter $\bm{\theta}^{*}\notin\bm{\Lambda}$, the interpolation of POD basis is needed to obtain $\bm{\Phi}^{*}$.
Grassmann Manifold. In honour of Hermann Grassmann, a Grassmann Manifold, $\mathcal{G}(r,n):=\{\mathbb{W}\subset\mathbb{R}^{n},\mathrm{dim}(\mathbb{W})=r\}$, is a differentiable manifold that collects all $r$-dimensional linear subspaces in $\mathbb{R}^{n}$[5]. Hence, the subspace $\mathbb{W}$ is a point in $\mathcal{G}(r,n)$ and the optimization problem in (2) essentially seeks the optimal point in $\mathcal{G}(r,n)$ that minimizes the distance $\sum_{i=1}^{n_{T}}||\bm{x}(t_{i};\bm{\theta})-\mathcal{W}(\bm{x}(t_{i};\bm{\theta}))||^{2}$.
Stiefel manifold and a submersion. A subspace in $\mathcal{G}(r,n)$ is spanned by an orthogonal basis $\bm{\Phi}\in\mathbb{R}^{n\times r}$ ($\bm{\Phi}^{T}\bm{\Phi}=\bm{I}_{r}$). A collection of $n\times r$ matrices with orthogonal columns forms a compact Stiefel manifold denoted by $\mathcal{ST}(r,n)$[12]. Note that, if $\bm{\Phi}$ spans a subspace $p\in\mathcal{G}(r,n)$, so does $\bm{\Phi}\bm{J}$ for any $r\times r$ matrix $\bm{J}$ such that $\bm{J}^{T}\bm{J}=\bm{I}_{p}$. Therefore, there exists is a fiber bundle
(3) $\pi:\bm{\Phi}\in\mathcal{ST}(r,n)\mapsto\pi(\bm{\Phi})=p\in\mathcal{G}(r,n)$
which indicates that any point $p\in\mathcal{G}(r,n)$ can be represented by a point of $\pi^{-1}(p)$.
Riemannian metric. To establish the Riemannian metric on $\mathcal{G}(r,n)$, a unique tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$ is attached to a point $p\in\mathcal{G}(r,n)$. The tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$ has the same dimension of $\mathcal{G}(r,n)$ (but is isomorphic to $\mathbb{R}^{r\times(n-r)}$, i.e., $\mathcal{T}_{p}\mathcal{G}(r,n)\simeq\mathbb{R}^{r\times(n-r)}$) and is equipped with a scalar product. For any two velocity vectors, $v_{1},v_{2}\in\mathcal{T}_{p}\mathcal{G}(r,n)$, we have
(4) $\left\langle v_{1},v_{2}\right\rangle:=\left\langle\bm{Z}_{1},\bm{Z}_{2}\right\rangle=\mathrm{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$
where $\bm{Z}_{1}$ and $\bm{Z}_{2}$ are $n\times r$ matrices from the horizontal space of $\bm{\Phi}=\pi^{-1}(p)$ denoted by $\mathbb{H}_{\bm{\Phi}}:=\{\bm{Z}\in\mathbb{R}^{n\times r};\bm{Z}^{T}\bm{\Phi}=0\}$. In other words, given a base point $p$, the Grassmann Manifold inherits the Riemannian structure of the Stiefel manifold due to (3) and becomes a Riemannian manifold.
Exponential and Logarithm maps. For any point $p\in\mathcal{G}(r,n)$ and the attached tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, the Exponential map is defined as
(5) $\text{Exp}_{p}:v\in\mathcal{T}_{p}\mathcal{G}(r,n)\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n)$
where $\bm{\Phi}=\pi^{-1}(p)$ and $\bm{Z}=\bm{U}_{\bm{Z}}\bm{\Sigma}_{\bm{Z}}\bm{V}_{\bm{Z}}$ is the SVD of $\bm{Z}$.
For any point $p\in\mathcal{G}(r,n)$ and a open set $U_{\bm{p}}:=\{p^{\prime}\in\mathcal{G}(r,n);\pi(\bm{\Phi})=p,\pi(\bm{\Phi}^{\prime})=p^{\prime},\pi(\bm{\Phi})^{T}\pi(\bm{\Phi}^{\prime})\text{ is invertible)}$), the Logarithm map is defined as
(6) $\text{Log}_{p}:p^{\prime}\in U_{\bm{p}}\mapsto\text{Log}_{p}(p^{\prime})\in\mathcal{T}_{p}\mathcal{G}(r,n).$
Existing POD interpolation method. Within $U_{\bm{p}}$, the Logarithm map $\text{Log}_{p}$ is a diffemorphism, and the one-one correspondence between $v\in\mathcal{T}_{p}\mathcal{G}$ and $p^{\prime}\in U_{\bm{p}}$ can be defined through the Exponential and Logarithm maps[12]. Hence, [2] described a well-known Lagrangian interpolation method for POD basis. In a nutshell, given the training data $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, subspaces corresponding to $\bm{\theta}_{i}$ are mapped to vectors $v_{i}$ in the tangent space through the Logarithm map. Then, the interpolation of vectors can be done in the tangent space, and the interpolated vector $v^{*}$ is mapped back to the Grassmann Manifold through the Exponential map. Such a POD basis interpolation algorithm is provided in Algorithm 1.
Algorithm 1 POD basis interpolation on Grassmann Manifold
  Input: Integers $r$ and $n$ ($2r\leq n$), $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, and a target parameter $\bm{\theta}^{*}$.
  • Choose a matrix $\bm{\Phi}_{0}\in\{\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}\}$ such that $\bm{\Phi}_{0}^{T}\bm{\Phi}_{i}$ is non-singular for all $i$.
  • For each $i=1,2,\cdots,k$, perform the thin SVD on $\bm{\Phi}_{i}(\bm{\Phi}_{0}^{T}\bm{\Phi}_{i})^{-1}-\bm{\Phi}_{0}=\bm{U}_{i}\bm{\Sigma}_{i}\bm{V}_{i}^{T}$, and obtain a matrix $\bm{Z}_{i}$
    (7) $\bm{Z}_{i}=\bm{U}_{i}\text{arctan}(\bm{\Sigma}_{i})\bm{V}_{i}^{T}.$
  • Compute an interpolated matrix and a thin SVD
    (8) $\bm{Z}^{*}=\sum_{i=1}^{k}\prod_{i\neq j}\frac{\bm{\theta}^{*}-\bm{\theta}_{j}}{\bm{\theta}_{i}-\bm{\theta}_{j}}\bm{Z}_{i}=\bm{U}^{*}\bm{\Sigma}^{*}\bm{V}^{*}$
  • Return an instability message if the largest singular value of $\bm{Z}^{*}$ is greater than $\pi/2$. Otherwise, compute a $n\times r$ matrix
    (9) $\bm{\Phi}^{*}=\bm{\Phi}_{0}\bm{V}^{*}\text{cos}\bm{\Sigma}^{*}+\bm{U}^{*}\text{sin}\bm{\Sigma}^{*}$
    where sin and cos only act on the diagonal of $\bm{\Sigma}^{*}$.
  return  The interpolated POD basis $\bm{\Phi}^{*}$ for parameter $\bm{\theta}^{*}$.
3 Projected Gaussian Process for Adapting ROMs
In this section, we describe a supervised statistical learning method, based on the proposed pGP, for adapting ROMs against parameter changes.
3.1 Projection from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$
We first show that there exists an injective mapping from the Euclidean space $\mathbb{R}^{nr-r}$ to the Grassmann Manifold $\mathcal{G}(r,n)$. This result will later enable us to define a GP on $\mathbb{R}^{nr-r}$ and map it to $\mathcal{G}(r,n)$.
From the preliminaries in Sec.2, there is an isomorphism between the horizontal space $\mathbb{H}_{\bm{\Phi}}$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, i.e., $\pi_{\mathbb{H}}:\mathbb{H}_{\bm{\Phi}}\rightarrow\mathcal{T}_{p}\mathcal{G}(r,n)$. Hence, it is possible to find a unique $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$ such that $\pi_{\mathbb{H}}(\bm{Z})=v\in\mathcal{T}_{p}\mathcal{G}(r,n)$, and the matrix $\bm{Z}$ is called the horizontal lift of the vector $v\in\mathcal{T}_{p}\mathcal{G}(r,n)$. The existence of such an isomorphism may suggest that a matrix-variate GP can be employed to model the relationship between $\bm{\theta}$ and the matrix $\bm{Z}$. However, it is important to point out that this seemingly natural approach needs to be carefully carried out because the matrix $\bm{Z}$ should be defined on the horizontal space $\mathbb{H}_{\bm{\Phi}}$ given a base point $p$, rather than the Euclidean space $\mathbb{R}^{n\times r}$. Precisely speaking, if $\bm{Z}$ is modeled as a GP with respect to the parameter $\bm{\theta}$ on $\mathbb{R}^{n\times r}$, then, the point prediction, $\hat{\bm{Z}}^{*}$, is given by a linear combination of $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ based on the well-known results of GP regression[28]. Since $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ are in the horizontal space $\mathbb{H}_{\bm{\Phi}}$ , the point prediction $\hat{\bm{Z}}^{*}$ remains in the same space. However, the uncertainty quantification or predictive internal of $\hat{\bm{Z}}^{*}$ (or, the predicted $\hat{\bm{\Phi}}^{*}$) is problematic because a sample $\bm{Z}$ from a Gaussian distribution on $\mathbb{R}^{n\times r}$ does not necessarily belong to $\mathbb{H}_{\bm{\Phi}}$.
To remedy this issue, we first present the following results:
Proposition 3.1.
Given a basepoint $p\in\mathcal{G}(r,n)$ and $\bm{\Phi}=[\bm{\phi}_{1},\cdots,\bm{\phi}_{r}]$ such that $\pi(\bm{\Phi})=p$, let $\bm{y}$ be a $(nr-r)\times 1$ vector, $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ be a $r\times nr$ matrix with orthogonal rows, and $\bm{F}=[\bm{F}_{1}|\bm{F}_{2}|\cdots|\bm{F}_{r}]$ be a $(nr-r)\times nr$ matrix where each block $\bm{F}_{i}$, $i=1,2,\cdots,r$, is a $(nr-r)\times n$ matrix that satisfies the following conditions: (i) for any block $i$ ($i=1,2,\cdots,r$), only row $(i-1)\times(n-1)+1$ to row $i\times(n-1)$ contain non-zero entries; (ii) all non-zero row vectors are unit vectors; and (iii) all non-zero row vectors are orthogonal to each other (note that conditions (i)-(iii) imply $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$). Then,
(10) $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{T}\bm{y})$
is a $n\times r$ matrix that belongs to the horizontal space $\mathbb{H}_{\bm{\Phi}}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$.
Proof. Note that $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$, i.e.,
(11) $\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=0$
where $\text{vec}(\bm{Z})$ is the column-wise vectorization of $\bm{Z}$, and $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ is a $r\times nr$ block diagonal matrix with orthogonal rows. Then, by introducing a $(nr-r)\times nr$ matrix $\bm{F}$ with orthogonal rows such that $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$, we have
(12) $\begin{pmatrix}\tilde{\bm{\Phi}}\\ \bm{F}\end{pmatrix}\text{vec}(\bm{Z})=\begin{pmatrix}\bm{0}\\ \bm{y}\end{pmatrix}$
where $\bm{y}$ is a $(nr-r)$-dimensional column vector.
Because $\bm{F}\text{vec}(\bm{Z})=\bm{y}$ is an underdetermined system, we may find $\text{vec}(\bm{Z})$ by $\bm{F}^{\dagger}\bm{y}$ where $\bm{F}^{\dagger}=\bm{F}^{T}(\bm{F}\bm{F}^{T})^{-1}$. From conditions (i)-(iii) above, $\bm{F}\bm{F}^{T}$ is an identity matrix, and $\bm{F}^{\dagger}\bm{F}^{T}$ as shown in Proposition 3.1.
Here, the choice of $\bm{F}$ depends on $\bm{\Phi}$ (and the basepoint $\bm{p}$ in Proposition 3.1) as it must satisfy conditions (i)-(iii). Let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}_{1},\tilde{\bm{\Phi}}_{2},\cdots,\tilde{\bm{\Phi}}_{r}]$, and let $\tilde{\bm{F}}=[\tilde{\bm{F}}_{1},\tilde{\bm{F}}_{2},\cdots,\tilde{\bm{F}}_{r}]$ be a $(nr-r)\times nr$ proposal matrix where each block $\tilde{\bm{F}}_{i}$ has the structure as specified by condition (i) above. Then, applying the Gram-Schmidt (G-S) process can be applied to the row vectors of each block $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \tilde{\bm{F}}_{i}\end{smallmatrix}\bigr{)}$ (for $i=1,2,\cdots,r$) and normalizing each non-zero rows after the G-S process, we obtain $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \bm{F}_{i}\end{smallmatrix}\bigr{)}$ where $\bm{F}_{i}$ satisfies conditions (i)-(iii). Note that, the G-S process starts from the $i$-th row because the leading $(i-1)$ rows of $\tilde{\bm{\Phi}}_{i}$ are zeros.
Proposition 3.1 implies that, from a vector $\bm{y}\in\mathbb{R}^{nr-r}$, it is possible to compute a matrix $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ that comes from the horizontal space $\mathbb{H}_{\bm{\Phi}}$. In fact, it is not hard to understand why the dimension of $\bm{y}$ is $nr-r$. Although there are $n\times r$ elements in the matrix $\bm{Z}$, the constraint $\bm{Z}^{T}\bm{\Phi}=\bm{0}$ (i.e., $\bm{Z}$ must come from the horizontal space of $\bm{\Phi}$) takes $r$ degree of freedom away, leaving behind $nr-r$ degrees of freedom. Hence, for any basepoint $p\in\mathcal{G}(r,n)$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, we can define the following mapping from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$,
(13) $\mathcal{P}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n).$
Here, $\mathcal{P}_{p}:=\tilde{\mathcal{P}}_{p}\circ\text{Exp}_{p}$ where $\tilde{\mathcal{P}}_{p}$ maps an $(nr-r)$-dimensional vector to the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$,
(14) $\tilde{\mathcal{P}}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto v\in\mathcal{T}_{p}\mathcal{G}(r,n)$
and the Exponential map, $\text{Exp}_{p}$, is defined in (5).
3.2 Injectivity of the map, $\mathcal{P}_{p}$
Although $\mathcal{P}_{p}$ in (13) defines a map from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$, it is important to guarantee that such a map is injective (i.e., a subspace in $\mathcal{G}(r,n)$ can be uniquely determined by a vector in $\Omega_{p}$). Proposition 3.2 below gives the condition for which $\mathcal{P}_{p}$ is injective.
Proposition 3.2.
Given a basepoint $p\in\mathcal{G}(r,n)$, there exists an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$
(15) $\Omega_{p}:=\left\{\bm{y}\in\mathbb{R}^{nr-r};||\bm{y}||<\frac{\pi}{2}\right\}$
such that the mapping $\mathcal{P}_{p}$ is injective.
Proof. Consider an open set $D_{p}\subset\mathcal{T}_{p}\mathcal{G}(r,n)$
(16) $D_{p}:=\left\{v\in\mathcal{T}_{p}\mathcal{G}(r,n);||v||<\pi/2\right\}$
where $\pi/2$ is the injective radius on the Grassmann Manifold[12]. Hence, we seek an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$ such that $\mathcal{A}_{p}$ maps a vector $\bm{y}\in\Omega_{p}$ to the open disk $D_{p}$ above. Because $\bm{Z}=\text{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$, we have
(17) $\begin{split}\bm{Z}^{T}\bm{Z}&=\begin{bmatrix}\bm{y}^{T}\bm{F}_{1}\\ \bm{y}^{T}\bm{F}_{2}\\ \vdots\\ \bm{y}^{T}\bm{F}_{r}\end{bmatrix}\begin{bmatrix}\bm{F}_{1}^{T}\bm{y}&\bm{F}_{2}^{T}\bm{y}&\cdots&\bm{F}_{r}^{T}\bm{y}\end{bmatrix}\\ &=\begin{bmatrix}\sum_{i=1}^{n-1}y_{(i-1)\times r+1}^{2}&&&\\ &\sum_{i=1}^{n-1}y_{(i-1)\times r+2}^{2}&&\\ &&\ddots&\\ &&&\sum_{i=1}^{n-1}y_{(i-1)\times r+r}^{2}\end{bmatrix}\end{split}$
Because the Riemannian metric on $\mathcal{G}(r,n)$ is defined by $\left\langle v_{1},v_{2}\right\rangle_{p}=\text{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$ given the basepoint $p$, we have
(18) $||v||=\sqrt{\text{trace}(\bm{Z}^{T}\bm{Z})}=\left(\sum_{i=1}^{nr-r}y_{i}^{2}\right)^{\frac{1}{2}}=||\bm{y}||$
which needs to be smaller than $\pi/2$. Hence, we have an open set $\Omega_{p}$, defined in (15), such that $\mathcal{P}_{p}$ is injective.
3.3 Prediction based on the pGP regression
Proposition 3.1 suggests an effective approach to establish the mapping between the parameter and POD basis. In a nutshell, as illustrated in Figure 3, a GP can be established to describe the mapping between the parameters space $\mathbb{R}^{d}$ and $\mathbb{R}^{nr-r}$, and then, the GP is projected to the Grassmann Manifold. As a result, for any new parameter $\bm{\theta}^{*}$, a vector $\bm{y}^{*}$ can be found and mapped back to the Grassmann Manifold to determine the optimal subspace onto which the full-order system is projected (equivalently, the POD basis matrix $\bm{\Phi}^{*}$).
Figure 3: A sketch of the key relationships between different spaces/manifolds involved in the pGP regression.
Let the vector $\bm{y}$ follows the multivariate Gaussian distribution, i.e., $\bm{y}\sim\mathcal{N}(\bm{0}_{nr-r,1},\bm{K})$ where $\bm{K}$ is a $(nr-r)\times(nr-r)$ symmetric and positive semi-definite matrix. Then, $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ follows a matrix-variate Gaussian, $\bm{Z}\sim\mathcal{N}(\bm{0}_{n\times r},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$ in the sense that $\text{vec}(\bm{Z})\sim\mathcal{N}(\bm{0}_{nr\times 1},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$. For a given point $\mu\in\mathcal{G}(r,n)$ and a symmetric positive semi-definite matrix $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, we say that a random point $X\in\mathcal{G}(r,n)$ follows a Projected Gaussian Distribution (PGD) on $\mathcal{G}(r,n)$
(19) $X\sim\mathcal{P}_{\mu}(\mathcal{N}(\bm{0}_{nr-r,1},\bm{K}))$
where the projection, $\mathcal{P}_{\mu}:\mathbb{R}^{nr-r}\rightarrow\mathcal{G}(r,n)$, is associated with the given point $\mu$. We denote the PGD by $\mathcal{N}_{\mathcal{G}}(\mu,\bm{K})$, and write $X\sim\mathcal{N}_{\mathcal{G}}(\mu,\bm{K})$, $\mu:=\mu_{\mathcal{G}}(X)\in\mathcal{G}(r,n)$, and $\text{cov}_{\mathcal{G}}(X):=\bm{K}$.
To obtain the mapping $\mathcal{P}_{\mu}$ that connects $\mathbb{R}^{nr-r}$ and $\mathcal{G}(r,n)$ given a point $\mu\in\mathcal{G}(r,n)$ and $\bm{\Phi}=\pi^{-1}(\mu)\in\mathcal{ST}(r,n)$, we first introduce an operator $\mathcal{P}_{1}$ such that $\mathcal{P}_{1}(\bm{y})=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})=\bm{Z}$. Then, let $\bm{Z}=\bm{U}_{\bm{Z}}\bm{\Sigma}_{\bm{Z}}\bm{V}_{\bm{Z}}^{T}$ be the thin SVD, an operator $\mathcal{P}_{2}$ can be introduced that operates on the SVD of $\bm{Z}$ such that $\mathcal{P}_{2}(\bm{Z})=\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}})$ yields a $n\times r$ matrix that spans a subspace in $\mathcal{G}(r,n)$. Hence, the mapping $\mathcal{P}_{\mu}$ can be defined as
(20) $\mathcal{P}_{\mu}:=\text{span}\circ\mathcal{P}_{2}\circ\mathcal{P}_{1}$
that maps a $(nr-r)\times 1$ vector from $\mathbb{R}^{nr-r}$ to a $r$-dimensional subspace in $\mathcal{G}(r,n)$.
Proposition 3.3.
If $X_{i}\sim\mathcal{N}_{\mathcal{G}_{i}}(\mu_{i},\bm{K}_{i})$ for $i=1,2$, then, $(X_{1},X_{2})$ are jointly PGD on $\mathcal{G}_{1}\times\mathcal{G}_{2}$
(21) $(X_{1},X_{2})\sim\mathcal{N}_{\mathcal{G}_{1}\times\mathcal{G}_{2}}\left(\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix},\tilde{\bm{K}}\right),\quad\quad\tilde{\bm{K}}=\begin{pmatrix}\bm{K}_{1},\bm{K}_{12}\\ \bm{K}_{12}^{T},\bm{K}_{2}\end{pmatrix}$
Then, the conditional distribution of $X_{1}$ given $X_{2}=p_{2}$ is given by
(22) $X_{1}\mid(X_{2}=p_{2})\sim\mathcal{P}_{\mu_{1}}(\mathcal{MN}(\bm{u}_{\bm{s}},\bm{K}_{\bm{s}}))$
where $\bm{u}_{\bm{s}}=\bm{K}_{12}\bm{K}_{2}^{-1}\bm{s}$, $\bm{K}_{\bm{s}}=\bm{K}_{1}-\bm{K}_{12}\bm{K}_{2}^{-1}\bm{K}_{12}^{T}$, $\bm{s}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$, and $\mathcal{P}_{\mu_{2}}^{-1}$ is the inverse map.
Proposition 3.3 is obtained following a similar technique in [21]. Let $\mathbb{A}_{1}=\mathcal{P}_{\mu_{1}}^{-1}(p_{1})=\{\bm{y}\in\mathbb{R}^{nr-r}:\mathcal{P}_{\mu_{1}}(\bm{Y})=p_{1}\}$ be the preimage of $p_{1}$ in $\mathbb{R}^{nr-r}$ associated with $\bm{\Phi}_{\mu_{1}}$ given a point $\mu_{1}$, and let $\mathbb{A}_{2}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$ similarly be the preimage of $p_{2}$, then,
(23) $\begin{split}\text{Pr}(X_{1}|X_{2}=p_{2})&=\frac{\text{Pr}(\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2})}{\text{Pr}(\bm{y}_{2}\in\mathbb{A}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\frac{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}{\text{Pr}(\mathbb{A}_{2})}\frac{\mathcal{MN}((\bm{y}_{1}^{T},\bm{y}_{2}^{T})^{T};\bm{0},\tilde{\bm{K}})}{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\lambda_{\bm{y}_{2}}\mathcal{MN}(\bm{y}_{1};\bm{u}_{\bm{y}_{2}},\bm{K}_{\bm{y}_{2}})\end{split}$
where $\bm{u}_{\bm{y}_{2}}=\bm{K}_{12}\bm{K}_{2}^{-1}\bm{y}_{2}$, and $\bm{K}_{\bm{y}_{2}}=\bm{K}_{1}-\bm{K}_{12}\bm{K}_{2}^{-1}\bm{K}_{12}^{T}$.
Hence,
(24) $X_{1}\mid(X_{2}=p_{2})\sim\mathcal{P}_{\mu_{1}}(\mathcal{MN}(\bm{u}_{\bm{s}},\bm{K}_{\bm{s}}))$
where $\bm{s}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$.
Once a PGD is obtained as described above, a Projected Gaussian Process (pGP) on Grassmann Manifold can be defined. We start with a multivariate GP defined in the Euclidean space. A collection $\bm{y}$ of random vectors is a Multivariate GP (MGP) in $\mathbb{R}^{d}$, denoted by $\bm{y}\sim\mathcal{MGP}(f,\omega,\bm{K})$, with a matrix-valued mean function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{nr-r}$, kernel $\omega:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, if any finite collection of $\bm{y}_{1}$, $\bm{y}_{2}$, $\cdots$, $\bm{y}_{k}$ have a joint multivariate Gaussian distribution such that
(25) $[\bm{y}_{1}^{T},\bm{y}_{2}^{T},\cdot,\bm{y}_{k}^{T}]^{T}\sim\mathcal{N}([f^{T}(\bm{\theta}_{1}),f^{T}(\bm{\theta}_{2}),\cdot,f^{T}(\bm{\theta}_{k})^{T}]^{T},\bm{\Omega}\otimes\bm{K})$
where $\bm{\Omega}\in\mathbb{R}^{k\times k}$ with $\bm{\Omega}_{i,j}=\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$. Then, a pGP on $\mathcal{G}(r,n)$ can be defined as follows:
Projected Gaussian Process. A Projected Gaussian Process (pGP) on $\mathcal{G}(r,n)$ is defined by
(26) $g\sim\mathcal{PGP}(m,\omega,\bm{K})$
where a collection $(g(\bm{\theta}_{1}),g(\bm{\theta}_{2}),\cdots,g(\bm{\theta}_{k}))\sim\mathcal{P}_{m}(\mathcal{MGP}(\bm{0},\omega,\bm{K})$ (i.e., a joint PGD), $m(\bm{\theta}):=\mu_{\mathcal{G}}(g(\theta))$ is the basepoint function and $k(\bm{\theta}_{i},\bm{\theta}_{i}):=\text{cov}_{\mathcal{G}}(g(\bm{\theta}_{i}),g(\bm{\theta}_{j}))=\omega(\bm{\theta}_{i},\bm{\theta}_{j})\bm{K}$ is the co-variance function in $\mathbb{R}^{nr-r}$. We denote the process by $g\sim\mathcal{PGP}(m,\omega,\bm{K})$.
Following Proposition 3.3, the joint distribution between the training output $\bm{p}^{\text{train}}=(p_{1},\cdots,p_{k})^{T}$ and test output $p_{*}$ is given by:
(27) $\begin{pmatrix}p_{*}\\ \bm{p}^{\text{train}}\end{pmatrix}\sim\mathcal{N}_{\mathcal{G}_{*}\times\mathcal{G}_{1}\times\cdots\times\mathcal{G}_{k}}\left(\begin{pmatrix}m_{*}\\ \bm{m}^{\text{train}}\end{pmatrix},\begin{pmatrix}\bm{K}_{**},\bm{K}_{*}\\ \bm{K}_{*}^{T},\bm{K}^{\text{train}}\end{pmatrix}\right)$
where $m_{*}=m(\theta_{*})$, $\bm{m}^{\text{train}}=(m(\bm{\theta}_{1}),m(\bm{\theta}_{2}),\cdots,m(\bm{\theta}_{d}))^{T}$, $\bm{K}^{\text{train}}=\bm{\Omega}\otimes\bm{K}\in\mathbb{R}^{k(nr-r)\times k(nr-r)}$, $K_{**}=\omega(\bm{\theta}_{*},\bm{\theta}_{*})\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, $\bm{K}_{*}=(\omega(\bm{\theta}_{*},\bm{\theta}_{1}),\omega(\bm{\theta}_{*},\bm{\theta}_{2}),\cdots,\omega(\bm{\theta}_{*},\bm{\theta}_{d}))\otimes\bm{K}\in\mathbb{R}^{(nr-r)\times k(nr-r)}$.
Finally,
(28) $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}))$
where
(29) $\bm{u}_{*}=\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}((\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T},(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T},\cdots,(\mathcal{P}_{m(\bm{\theta}_{d})}^{-1}(p_{k}))^{T})^{T}$
and
(30) $\tilde{\bm{K}}_{*}=\bm{K}_{**}-\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}\bm{K}_{*}^{T}.$
Hence, a point estimate of the subspace $p_{*}\in\mathcal{G}(r,n)$ can be obtained by $\hat{p}_{*}=\mathcal{P}_{m_{*}}(\bm{u}_{*})$. Because $\bm{u}_{*}$ is the mean of the multivariate Gaussian distribution $\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*})$ defined in $\mathcal{R}^{nr-r}$, $\hat{p}_{*}$ is a maximum a posteriori probability (MAP) estimate. Also recall that $||\bm{u}_{*}||$ needs to be smaller than $\pi/2$ for $\mathcal{P}_{m_{*}}$ to be injective. Hence, if $||\bm{u}_{*}||>\pi/2$, we shrink this vector $\bm{u}_{*}\leftarrow\frac{\pi}{2}\frac{\bm{u}_{*}}{||\bm{u}_{*}||}$, which is the projection of the original vector $\bm{u}_{*}$ onto the surface of a $(nr-r)$-dimensional sphere. Similarly, the estimate $\hat{\bm{\Phi}}_{*}$ is obtained as $\mathcal{P}_{2}\circ\mathcal{P}_{1}(\bm{u}_{*})$ such that $\hat{p}_{*}$ is spanned by $\hat{\bm{\Phi}}_{*}$.
Figure 4: An illustration of the subspace prediction using the pGP regression.
Figure 4 illustrate the process of predicting the optimal subspace using pGP. In this illustration, suppose that we obtain the snapshot data from three parameter settings $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$, as well as the optimal subspaces $p_{1}$, $p_{2}$ and $p_{3}$ that can be found using POD. The prediction of optimal subspace involves the following steps:
(i) Given the mean function $m(\bm{\theta})$, the three subspaces $p_{1}$, $p_{2}$ and $p_{3}$ are firstly mapped to the Euclidean space $\mathbb{R}^{nr-r}$ through the projection $\mathcal{P}^{-1}_{m(\bm{\theta})}$; see (29). Note that, $(\mathcal{P}_{m(\bm{\theta})}^{-1}(p))^{T}$ returns a $(nr-r)\times 1$ column vector on $\mathbb{R}^{nr-r}$.
(ii) The three column vectors $(\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T}$, $(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T}$ and $(\mathcal{P}_{m(\bm{\theta}_{3})}^{-1}(p_{3}))^{T}$ are used to predict a new column vector $\bm{u}_{*}$, for the new parameter $\bm{\theta}_{*}$, through a linear combination defined by a GP; see (29).
(iii) The covariance matrix $\tilde{\bm{K}}_{*}$ is computed on $\mathbb{R}^{nr-r}$, which is exactly the same as the convectional GP. Steps (ii) and (iii) together yield a multivariate Gaussian distribution of $\bm{y}_{*}$) on $\mathbb{R}^{nr-r}$, i.e., $\bm{y}_{*}\sim\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*})$.
(iv) Finally, the Gaussian distribution of $\bm{y}_{*}$ on $\mathbb{R}^{nr-r}$ (or, the predicted column vector $\bm{u}_{*}$ if only a point prediction is needed), is projected to the Grassmann Manifold through the projector $\mathcal{P}_{m(\bm{\theta}_{*})}$.
3.4 Discussions on model specification and parameter estimation
A close examination of (28) reveals an important difference between how prediction is done by conventional GP and the proposed pGP. Using a conventional GP, the point predictor is a linear combination of the observations and the weights are determined by the covariance function (i.e., the point prediction does not depend the mean of the GP; a property well-known in the literature). For the proposed pGP, as indicated by (28), the point prediction does depend on the function $m(\bm{\theta})$ that describes the relationship between the parameter and basepoint on the Grassmann Manifold. In other words, the choice of $m(\bm{\theta})$ determines the mapping $\mathcal{P}_{m}$. Two approaches can be used to determine the function $m(\bm{\theta})$. The first approach is to fix the basepoint for all parameters, i.e., $m(\bm{\theta})=\mu$. which can be obtained using a global POD basis. The second approach is to use another regression model to determine $m(\bm{\theta})$, such as the tree-based method described in [18].
In addition, because $n$ is usually large, a practical way to parameterize the $(nr-r)\times(nr-r)$ covariance matrix $\bm{K}$ is to let $\bm{K}=\sigma_{\bm{K}}\bm{I}$ and estimate $\sigma_{\bm{K}}$ from data. As for the kernel function $\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$, existing Gaussian, Exponential or other kernel functions can be used, and the parameter $\bm{\xi}$ can be estimated from data. Hence, once $m(\bm{\theta})$ is determined, the cross-validation-based approach can be used to estimate $(\sigma_{\bm{K}},\bm{\xi})$ by minimizing $\sum_{i=1}^{k}\delta_{i}$, where the cross validation error is given by $\delta_{i}=||\bm{D}(\bm{\theta}_{i})-\hat{\bm{\Phi}}_{-i}\hat{\bm{\Phi}}_{-i}^{T}\bm{D}(\bm{\theta}_{i})||_{F}^{2}$, and $\hat{\bm{\Phi}}_{-i}$ is the estimated $\bm{\Phi}_{i}$ excluding data $\bm{D}(\bm{\theta}_{i})$.
Also note that, for systems with high-dimensional parameter space (i.e., relatively large $d$), it is often the case that not all parameters are equally important in terms of how they affect the POD bases. In other words, the ROM can be sensitive to some parameters while independent of the others. Hence, when constructing the statistical learning model that establishes the mapping from parameter space to Grassmann Manifold, it is necessary to identify those relevant parameters. For the pGP model, one can achieve this goal using the Automatic Relevance Determination (ARD)[28]. To illustrate the idea, consider the squared Exponential kernel given by
(31) $\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}$
where $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{4}^{-2},\cdots,\xi_{d+2}^{-2}\}$. Here, the parameters $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ play a role of characteristic length-scales for parameter $\theta_{1},\theta_{2},\cdots,\theta_{d}$. If the estimated length-scales are very large, the corresponding parameters have little effects on the prediction and are practically removed from the model.
Finally, it is worth noting that the proposed pGP naturally involves the use of the global POD basis as its special case. If $\xi_{1}=0$ and the estimated $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ are all close to zero, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ becomes practically zero. In this case, (29) immediately implies that $\bm{u}_{*}=\bm{0}$ and $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}))$. If the global POD basis is chosen as the reference point (which spans a subspace in the Grassmann Manifold where the tangent space is attached to) and the basepoint function $\bm{m}$ is invariant with $\bm{\theta}$, then, any subspace, $p_{i}$, corresponding to the ROM at parameter $\bm{\theta}_{i}$ on the Grassmann Manifold can be seen as i.i.d. samples from a PGD, $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}))$. Hence, the point prediction, $\hat{p}_{*}$, is just the subspace spanned by the global POD basis.
4 Numerical Examples
Three numerical experiments are performed to demonstrate the proposed method: (i) flow around a circular cylinder; (ii) heat transfer in powder bed fusion additive manufacturing; and (iii) advection-diffusion processes.
4.1 Example I: Flow around a Cylinder
We first introduce the experiment setup and then present the experiment results.
Experiment Setup. In Example I, we consider the flow passing over a cylinder in a rectangle domain. The Lattice Boltzmann method is used to simulate the flow, i.e., to generate the training data for our model [22]. In particular, the Bhatnagar–Gross–Krook (BGK) approximation to the Navier-Stokes equation is:
(32) $\dot{f}+(\bm{v}\cdot\triangledown)f=-\tau(f-\bar{f})$
where $f(\bm{x})$ be the phase-space density of fluids at location $\bm{x}$, $\bm{v}$ is the velocity, $\tau$ is the timescale of which collisions happen, and $\bar{f}$ represents the equilibrium state. In (32), the left side $(\frac{\partial}{\partial t}f+(\bm{v}\cdot\triangledown))f$ represents the streaming of particles, while the right side $-\tau(f-\bar{f})$ represents the collision. The model assumes that the fluid locally relaxes to equilibrium over a characteristic timescale $\tau$ related to the kinematic viscosity. The larger the $\tau$ is, the larger is the kinematic viscosity.
Figure 5: Simulated flow velocity at times ($\tau=0.53$)
The fluid equation (32) is solved by the Lattice Boltzmann method using the Python code in [22]. As shown in Figure 5, we consider a $400\times 100$ rectangle domain. A solid cylinder is placed at location $(100,50)$ with a radius of 13. The velocity goes from left to right with a magnitude of $v=2.3$ per time step, and 5000 time steps are simulated. The initial condition follows a Gaussian distribution with mean zero and standard deviation 0.01. Reflective boundary conditions are considered for the cylinder, the top and the bottom boundaries, while the fluid is absorbed by the right boundary. Figure 5 shows the velocity of the flow, with $\tau=0.53$, at times 1, 2000, 4000 and 5000. The Von Karman vortices are visible at times 4000 and 5000.
Figures 6 and 7 show the 1st, 4th, 7th and 10th POD modes for $\tau=0.53$ and $\tau=0.7$, respectively. We see that the POD modes change as the parameter $\tau$ changes, suggesting the necessity of adapting POD basis.
Figure 6: The 1st, 4th, 7th and 10th POD modes for $\tau=0.53$
Figure 7: The 1st, 4th, 7th and 10th POD modes for $\tau=0.7$
To evaluate the accuracy of the predicted ROM through cross validation, we generate data for 33 values of $\tau$ ranging from 0.53 to 0.7 with a step size of 0.005. Then, the data are divided into a training and a testing sets. The training set includes the data generated from the following 11 parameter settings,
$\tau^{\text{train}}=\{0.53,0.545,0.56,0.575,0.59,0.605,0.62,0.635,0.65,0.665,0.68\},$
while the data generated from the remaining 22 parameter settings are used for testing.
In particular, we let $m(\tau)$ yields a constant basepoint which is the subspace spanned by the global POD basis. The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$. Note that, $\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}$ does not depend on $\sigma_{\bm{K}}$ when $\bm{K}=\sigma_{\bm{K}}\bm{I}$, and only $\xi_{1}$ and $\xi_{2}$ are to be estimated by minimizing the cross-validation error. We obtain $\hat{\xi}_{1}\approx 1$ and $\hat{\xi}_{2}\approx 0.5$.
Experiment Results. Figure 8 shows the out-of-sample testing errors for the proposed approach, the existing POD interpolation method, and the use of global POD bases. The errors are measured by the $\ell_{2}$ Frobenious norm defined under (2). Some important observations are summarized as follows:
$\bullet$ The proposed approach yields much smaller $\ell_{2}$ error than the existing POD interpolation method from testing cases #1 to #6, as well as from #19 to #22 (i.e., towards the two boundaries of the parameter space of $\tau$). This consist observation suggests that the existing Lagrangian interpolation may have some stability issues for parameters near the boundaries of the parameter space, as investigated in [12]. In addition, as discussed in the introduction section, the proposed approach allows us to optimally estimate the model parameters based on data, while the coefficients of the Lagrangian interpolation are independent from data. Overall, the proposed approach yields smaller $\ell_{2}$ error than the existing POD interpolation method for 21 out of the 22 testing cases (except for testing case #14 where the $\ell_{2}$ errors are almost identical).
$\bullet$ The benefit of interpolating POD basis over the use of global POD bases is also demonstrated. The proposed approach yields smaller $\ell_{2}$ errors than the use of the global POD bases for 17 out of the 22 testing cases.
Figure 8: Comparison of out-of-sample testing errors measured by the Frobenious norm
Figure 9: Standard deviation associated with the ROM at time 100.
As a statistical learning approach, the proposed approach is capable of quantifying the uncertainty of the output generated from the ROM. Figure 9 shows the bootstrapping standard deviation of the output of the ROM for testing cases #1, #8, #15 and #22. Within each iteration, a subspace is randomly sampled from the distribution $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}))$ obtained from (28), and an ROM is constructed on that sampled space. This procedure is repeated for 1000 times, and the standard deviation of the ROM output at time 100 is computed for each spatial location.
4.2 Experiment II: Powder Bed Fusion Additive Manufacturing
Powder Bed Fusion (PBF) with metals or polymers is one of the most common 3D printing technologies in industrial additive manufacturing (AM). PBF works by jointing powdered material point by point using energy sources such as lasers or electron beams. Thermal analysis is essential for product quality control which is key to spreading laser-based AM techniques [33] and manufacturing process design [17].
Figure 10: Illustration of problem setup, spatial domain and boundary conditions.
Experiment Setup. In this example, we focus on the transient heat transfer in the single-track L-PBF (laser PBF) process as shown in Figure 10. The vacuum chamber is preheated to a specific temperature and the spatial domain is a combination of the solid substrate and the powder layer. The transient heat transfer process is governed by a PDE:
(33) $\rho c_{p}\frac{\partial u(x,t)}{\partial t}-\nabla\cdot k\nabla u(x,t)=0,\quad(x,t)\in\Omega\times[0,t_{f}],$
where $\rho$ denotes the density, $c_{p}$ is the specific heat capacity, $k$ is the thermal conductivity of the material, $u(x,t)$ represents the desired temperature over a space-time domain $\Omega\times[0,t_{f}]$. The heat flux boundary condition in PBF manufacturing processes is described as
(34) $(-k\nabla u(x,t))\cdot\mathbf{n}=q_{h}+q_{c}+q_{r},\quad(x,t)\in\partial\Omega\times[0,t_{f}],$
where $\mathbf{n}$ is the (outward) unit normal vector to the surface, $q_{h}$, $q_{c}$, and $q_{r}$ are the heat flux due to the heat source, convection and radiation, respectively.
A moving Gaussian heat source considered as follows:
(35) $q_{h}=-\frac{2\alpha\tau}{\pi\gamma^{2}}\text{exp}(-\frac{2\|x-x^{c}(t)\|^{2}}{\gamma^{2}}),\quad x\in\partial\Omega_{t},$
where $\partial\Omega_{t}$ denotes the top surface of domain, $\alpha$ is the heat absorptivity, $\tau$ is the laser power, $\gamma$ is the effective laser (or electron) beam radius, $x^{c}(t)$ is the center location of the energy beam at time $t$.
The convection and radiation, for $x\in\partial\Omega_{t}\cup\partial\Omega_{s}$, are modeled as
(36) $q_{c}(x,t)=h(u(x,t)-u_{a}),\quad q_{r}(x,t)=\sigma_{s}\epsilon(u(x,t)^{4}-u_{a}^{4}),$
where $\partial\Omega_{s}$ denotes the side surfaces of domain, $h$ is the heat convection coefficient, $u_{a}$ is the ambient temperature, $\sigma_{s}$ is the Stefan-Boltzmann constant, $\epsilon$ is the emissivity of the material. The bottom surface of domain is defined with a Dirichlet boundary, $u(x,\cdot)=u_{b}$, where $x\in\partial\Omega_{b}$, $\partial\Omega_{b}$ denotes the bottom surface of domain and $u_{b}$ is the fixed temperature. Finally, the initial condition is defined as $u(\cdot,0)=u_{0}$ where $x\in\Omega$ and $U_{0}$ is the initial temperature. Note that, $\partial\Omega=\partial\Omega_{t}\cup\partial\Omega_{s}\cup\partial\Omega_{b}$.
The Finite Element method (FEM) is used to solve the governing equations. Multiplying both sides of (33) by a test function $\phi$ and integrating over $\Omega$ gives
[…]
The solution $u(x,t)$ and the test function are assumed to belong to Hilbert spaces. Unlike the original formulation that (33) holds for all points in $\Omega$, the weak formulation requires (37) to hold for all test functions in test function space. The Galerkin method assumes that the solution $u(x,t)$ belongs to the same Hilbert space as the test functions. According to Green’s first identity, (37) gives
[…]
Then, the discretization by Galerkin method is applied to the weak formulation and the approximated solution $u_{d}(x,t)$ is given by $u_{d}(x,t)=\sum_{i=1}^{N}u_{i}(t)\psi_{i}(x)$, where $N$ is the number of nodes. For every test function $\psi_{j}(x)$, we have
[…]
Figure 11: Temperature field at 600W laser power with the heat source moving at a speed of 5mm/s.
We generate data for 7 laser power levels (in Watt): $\{500,550,600,650,700,750,800\}$, using the Finite Element method. As an illustration, Fig.11 shows the temperature field at selected times. Since seven laser power levels are used to generate the data, a Leave-One-Out Cross-Validation (LOOCV) is performed to compare the model performance. Similar to Example 1, we let $r=10$ and $m(\tau)$ be a constant basepoint which is the subspace spanned by the global POD basis. The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$.
Experiment Results. Figure 12 shows LOOCV testing $\ell_{2}$ errors for the proposed approach, the existing POD interpolation method, and the use of global POD bases. It is important note that,
$\bullet$ The use of global POD basis and the proposed pGP approach yield better and very similar results.
$\bullet$ More importantly, the proposed pGP is able to provide an explanation to such an observation. In short, this is because the proposed pGP (i) involves the use of global POD basis as its special case (due to its parameter selection capabilities), and (ii) successfully identifies that the ROM, in this example, is not sensitive to the change of parameters, and the use of the global POD basis becomes appropriate.
To elaborate, the kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$ in this example. When LOOCV is performed, we note that the estimated values for $\xi_{2}$ ranges between 40 to 80, effectively making $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ close to zero. In this case, (28) immediately implies that $\bm{u}_{*}=\bm{0}$ and $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}))$. Because the global POD basis is chosen as the reference point, then, any subspace (corresponding different power levels $\tau$) on the Grassmann Manifold can be seen as i.i.d. samples from a PGD, $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}))$. Hence, the point prediction, $\hat{p}_{*}$, is close to the subspace spanned by the global POD basis, justifying the observation in Figure 12.
Figure 12: Leave-One-Out Cross-Validation (LOOCV) errors.
4.3 Experiment III: Advection-Diffusion Processes
Advection-diffusion equations are widely used to describe the transfer of particles, energy, or other physical quantities inside a physical system due to diffusion and advection.
Experiment Setup. We consider a general form of advection-diffusion process $\mathcal{A}({\bm{\theta}})x(t,\bm{s})=0$, where $x(t,\bm{s})$ is a spatio-temporal process in space $\bm{s}\in\mathbb{S}\subset\mathbb{R}^{2}$ and time $t\in[0,T]$, and $\mathcal{A}({\bm{\theta}})$ is the advection-diffusion operator parameterized by $\bm{\theta}$:
(40) $\mathcal{A}x:=\dot{x}+\bm{\vec{v}}^{T}\triangledown x-\triangledown\cdot[\bm{D}\triangledown x]$
[…]
Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}\in\mathbb{S}$, we represent the advection-diffusion process by a dynamical system that is linear in state
(41) $\bm{x}(t;\bm{\theta})=\bm{F}\bm{\alpha}(t;\bm{\theta}),\quad\dot{\bm{\alpha}}(t;\bm{\theta})=\bm{G}(\bm{\theta})\bm{\alpha}(t;\bm{\theta})$
Here, $\bm{F}$ is an $n\times n$ matrix with its $i$-th row being $\bm{f}(\bm{s}_{i})=(f_{1}(\bm{s}_{i}),...,f_{n}(\bm{s}_{i}))$ where $f_{j}(\bm{s})=\exp(i\bm{k}_{j}^{T}\bm{s})$ is the Fourier basis function and $\bm{k}_{j}$ is the spatial wavenumber, $\bm{\alpha}(t)\in\mathbb{R}^{n}$ is the state vector, $\bm{G}(\bm{\theta})$ is an $n\times n$ state transition matrix operator, and the initial condition is denoted by $\bm{x}_{0}$. Note that, the matrix $\bm{G}(\bm{\theta})$ depends on the parameters in (40) and is given by [19].
In projection-based ROM of parametric dynamic systems[6], we seek a ROM of the form to replace the original large-scale model (41) :
(42) $\bm{x}_{r}(t;\bm{\theta})=\bm{F}_{r}\bm{\alpha}_{r}(t;\bm{\theta}),\quad\dot{\bm{\alpha}}_{r}(t;\bm{\theta})=\bm{G}_{r}(\bm{\theta})\bm{\alpha}_{r}(t;\bm{\theta})$
where $\bm{\alpha}_{r}(t)$ is the $r$-dimensional reduced-order state vector ($r\ll n$), $\bm{x}_{r}(t)$ is an approximation of $\bm{\alpha}_{r}(t)$, and $\bm{F}_{r}\in\mathbb{R}^{n\times r}$ and $\bm{G}_{r}\in\mathbb{R}^{r\times r}$. Following the Galerkin’s method, it is known that $\bm{\alpha}_{r}(t)=\bm{\Phi}_{\bm{\alpha}}^{T}\bm{\alpha}(t)$, $\bm{G}_{r}=\bm{\Phi}_{\bm{\alpha}}^{T}\bm{G}\bm{\Phi}_{\bm{\alpha}}$ and $\bm{F}_{r}=\bm{F}\bm{\Phi}_{\bm{\alpha}}$, where $\bm{\Phi}_{\bm{\alpha}}$ is a matrix basis with orthogonal columns. Given any parameter $\bm{\theta}$, we obtain from the snapshot data matrix $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$ the POD basis $\bm{\Phi}$. Because $\bm{x}(t;\bm{\theta})=\bm{F}\bm{\alpha}(t;\bm{\theta})$, i.e., $\bm{x}(t;\bm{\theta})$ is the left unitary transformation of $\bm{\alpha}(t;\bm{\theta})$, the invariance property of SVD to unitary transformation yields $\bm{\Phi}_{\bm{\alpha}}=\bm{F}{\bm{\Phi}}$[16].
To evaluate the accuracy of adapting ROM using the proposed pGP, we let $\bm{v}=(v_{x},0.005)^{T}$, $\bm{D}=\mathrm{diag}\{0.001,0.001\}$ and generate data by letting $v_{x}$ vary from -0.02 to 0.02 with a step size of 0.001. The same initial condition is used for all $v_{x}$ values. In total, a number of 42 snapshot data matrices are generated for 42 values of $v_{x}$. The data are divided into a training and a testing sets. The training set includes the data generated from the following 11 parameter settings, $v_{x}^{\text{train}}=\{-0.016,-0.012,-0.008,-0.004,-0.002,0,0.004,0.008,0.012,0.016,0.02\}$, while the data generated from the remaining 31 parameter settings are used for testing. In the testing phase, the trained pGP model is used to predict the optimal POD basis $\bm{\Phi}_{\bm{\alpha}}$ for a given value of $v_{x}$, and the corresponding ROM (42) is constructed. Then, we solve the ROM to obtain the reduced-order solution $\bm{x}_{r}(t)$ and re-construct the full-order solutions by $\bm{x}(t)=\bm{\Phi}\bm{x}_{r}(t)$.
Experiment Results. To show and compare the performance of different approaches, Figure 13 shows the (re-constructed) solution from the proposed pGP, (re-constructed) solution from the existing POD interpolation approach, (re-constructed) solution using the global POD, and the (ground truth) solution by directly solving the full-order system. For illustrative purposes, these figures show the results for $v_{x}=0.006$ (which is not included in the training dataset) at times 1 (the initial condition), 5, 10, and 15. We see that the accuracy of the (reconstructed) solutions obtained from the existing POD interpolation approach deteriorates starting from time 5, while the proposed pGP approach provides more accurate solutions.
(a) t=1 (initial condition)
(b) t=5
(c) t=10
(d) t=15
Figure 13: Comparison between the proposed pGP (row 1 of each subfigure), the existing POD interpolation approach (row 2), the use of global POD (row 3), and the (ground truth) solution by directly solving the full-order system (row 4) at $v_{x}=0.006$
To further show the advantages of the proposed pGP approach, Figure 14 presents the comparison of the overall accuracy for all testing cases. The improvement by the proposed pGP is clearly observed.
(a) $\ell_{2}$ error
(b) $L_{\infty}$ error
Figure 14: Comparison between the proposed approach, existing POD interpolation approach and the use of global POD for constructing ROM against parameter changes.
4.4 Code and Data
Code and data are submitted as Supplementary Materials. Because the size of the dataset is about 3.5GB, users may download the data from the following link: https://drive.google.com/file/d/1Ub_TnaB0--cW3tcNlSsr2qiVdliH2xdJ/view?usp=sharing.
5 Conclusions
This work successfully demonstrated the great potential of introducing machine learning to the field of robust reduced-order modeling. A supervised learning problem, based on the proposed pGP regression, has been developed to predict the optimal POD basis. The proposed approach leverages an injective mapping from the parameter space to the Grassmann Manifold that contains the optimal vector subspaces. The condition for such a mapping to be injective has been obtained, and sufficient technical details were presented for predicting POD basis using the proposed pGP regression. Numerical experiments showed that the proposed pGP approach not only improves the accuracy of ROMs against parameter change, but also enables uncertainty quantification and parameter selection, potentially reducing the dimension of the parametric space when adapting ROMs. We would also like to point out that there exists a critical link between the proposed pGP and statistical experimental design/reinforcement learning that enable sequential exploration of a system’s behavior in a reduced parameter space. This link is expected to create important future research trajectories, potentially attracting more researchers into this field and making tangible contributions to Scientific Machine Learning research and its applications.

### v2

@@
-A Statistical Machine Learning Approach for Adapting Reduced-Order Models using Projected Gaussian Process
+Adapting Projection-Based Reduced-Order Models using Projected Gaussian Process
 Abstract
-Projection-based model reduction is among the most widely adopted approaches for constructing parametric Reduced-Order Models (ROMs).
-Utilizing snapshot data from solving full-order governing equations, the Proper Orthogonal Decomposition (POD) computes the optimal basis modes that span a low-dimensional subspace where the ROM resides.
-Because a governing equation is often parameterized by a set of parameters, challenges immediately arise when one would like to investigate how systems behave differently over the parameter space (in design, diagnosis, control, uncertainty quantification and real-time operations).
-In this case, the optimal basis needs to be efficiently updated so as to adapt ROM that can accurately capture the variation of a system’s behavior over its parameter space.
-In this paper, we propose a Projected Gaussian Process (pGP) model and formulate the problem of adapting POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping (injective) from the parameter space to the Grassmann Manifold that contains the optimal vector subspaces.
-To establish such a relationship, a mapping is found between the Euclidean space and the horizontal space of an orthogonal matrix that spans a reference subspace in the Grassmann Manifold.
-Then, a second mapping from the horizontal space to the Grassmann Manifold is established through the Exponential/Logarithm maps between the manifold and its tangent space.
-Finally, given a new parameter, the conditional distribution of a vector can be found in the Euclidean space using the Gaussian Process (GP) regression, and such a distribution is then projected to the Grassmann Manifold that enables us to find the optimal subspace, i.e., POD basis, for the new parameter.
-Compared with existing interpolation method, the proposed statistical learning approach allows us to optimally estimate (or tune) model parameters given data (i.e., the prediction/interpolation becomes problem-specific), and quantify the uncertainty associated with the prediction.
-Numerical examples are presented to demonstrate the advantages of the proposed pGP for adapting POD basis against parameter changes.
-keywords:
-Reduced-Order Models, Proper Orthogonal Decomposition, Gaussian Process, Projected Gaussian Process, Grassmann Manifold
+Projection-based model reduction is among the most widely adopted methods for constructing parametric Reduced-Order Models (ROM).
+Utilizing the snapshot data from solving full-order governing equations, the Proper Orthogonal Decomposition (POD) computes the optimal basis modes that represent the data, and a ROM can be constructed in the low-dimensional vector subspace spanned by the POD basis.
+For parametric governing equations, a potential challenge arises when there is a need to update the POD basis to adapt ROM that accurately capture the variation of a system’s behavior over its parameter space (in design, control, uncertainty quantification, digital twins applications, etc.).
+In this paper, we propose a Projected Gaussian Process (pGP) and formulate the problem of adapting the POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping from the parameter space to the Grassmann manifold that contains the optimal subspaces.
+A mapping is firstly established between the Euclidean space and the horizontal space of an orthogonal matrix that spans a reference subspace in the Grassmann manifold.
+A second mapping from the horizontal space to the Grassmann manifold is established through the Exponential/Logarithm maps between the manifold and its tangent space.
+Finally, given a new parameter, the conditional distribution of a vector can be found in the Euclidean space using the Gaussian Process (GP) regression, and such a distribution is then projected to the Grassmann manifold that enables us to predict the optimal subspace for the new parameter.
+As a statistical learning approach, the proposed pGP allows us to optimally estimate (or tune) the model parameters from data and quantify the statistical uncertainty associated with the prediction.
+The advantages of the proposed pGP are demonstrated by numerical experiments.
+Keywords: Reduced-Order Models, Grassmann manifold, Proper Orthogonal Decomposition, Gaussian Process, Projected Gaussian Process.
 1 Introduction
@@
 Consider the diffusion of chemical species, aeroelasticity of aircraft wings, flow of vehicles and predator-prey interactions, dynamics of such complex systems is typically described by a set of Partial Differential Equations (PDEs).
-Very often, the dimensions of these problems are extremely high due to the spatial discretization of the mathematical models (e.g., Computational Fluid Dynamics models or Finite Element Analysis with very large numbers of degrees of freedom).
-The prohibitive computational cost prevents us from repeatedly running numerical solvers under different parameter settings; often needed for design, diagnosis, prediction, uncertainty quantification and control problems[24, 25, 29].
-Hence, when there exist low-dimensional patterns embedded in high-dimensional systems[16], Reduced-Order Models (ROMs) become indispensable in capturing the dominate behaviors of complex systems.
-Compared with data-driven surrogate models or meta models[3], ROM often preserve important system physics, such as stability[26, 4], passivity[31, 9], polynomial nonlinearities[27], etc.
-Projection-based model reduction is among the most widely adopted approaches for constructing ROMs[16, 27].
-The central idea behind projection-based ROM is to project a high-dimensional governing equation to an optimal low-dimensional vector space spanned by orthogonal bases.
-Although searching for such bases may seem to be nontrivial considering the computational speed, accuracy and complex boundary conditions, the Proper Orthogonal Decomposition (POD)—applied to curves in Hilbert spaces $\mathcal{H}$ of infinite dimension—provides a computationally efficient method for approximating data generated from full-order systems in an $\ell_{2}$ sense[6].
-This can be efficiently done by performing the thin Singular Value Decomposition (SVD) of the snapshot data matrix[16].
-However, a governing equation is often parameterized by a set of parameters, and challenges immediately arise when one would like to investigate how systems behave differently over the parameter space (of the governing equation).
-In this case, the optimal basis needs to be updated so that the ROM can accurately capture the variation of a system’s behavior over its parameter space.
-As pointed out by [10, 32], the robustness and accuracy of an ROM are related to the reference parameters from which POD modes are derived.
-The accuracy of ROM determines the performance of any downstream machine learning models built on the low-dimensional ROM.
+Very often, the dimensions of these problems are high after the spatial discretization of the mathematical models (e.g., computational fluid dynamics models or finite element analysis with very large numbers of degrees of freedom).
+The prohibitive computational cost prevents us from repeatedly running numerical solvers under different parameter settings; often needed for design, diagnosis, prediction, uncertainty quantification and digital twins applications (Peherstorfer and Willcox, 2015a, ; Peherstorfer and Willcox, 2015b, ; Sargsyan et al.,, 2015).
+Hence, when there exist low-dimensional patterns embedded in high-dimensional systems (Kutz et al.,, 2016), Reduced-Order Models (ROM) become indispensable in capturing the dominate behaviors of complex systems.
+Compared with data-driven surrogate models or meta models (Barton and Meckesheimer,, 2006), ROM often preserve important system physics, such as stability (Prajna,, 2003; Batool et al.,, 2021), passivity (Sorensen,, 2003; Breiten and Unger,, 2022), polynomial nonlinearities (Qian et al.,, 2020), etc.
+Among various model reduction approaches, projection-based model reduction is perhaps the most widely adopted one for constructing ROM (Kutz et al.,, 2016; Qian et al.,, 2020).
+The central idea behind projection-based ROM is to project a high-dimensional system to an optimal low-dimensional vector space spanned by orthogonal bases.
+Although searching for such bases may seem to be a nontrivial task considering the computational speed, accuracy and complex boundary conditions, the Proper Orthogonal Decomposition (POD)—applied to curves in Hilbert spaces $\mathcal{H}$ of infinite dimension—provides a computationally efficient approach that optimally approximates data generated from full-order systems in an $\ell_{2}$ sense (Berkooz et al.,, 1993; Rowley,, 2005; Benner et al.,, 2015).
+This can be efficiently done by performing the thin Singular Value Decomposition (SVD) of the snapshot data matrix, and one may refer to Kutz et al., (2016) for a more detailed description of the procedure.
 To elaborate, consider a governing equation with parameters $\bm{\theta}=(\theta_{1},\theta_{2},...,\theta_{d})\in\mathbb{P}\subset\mathbb{R}^{d}$, where $\mathbb{P}$ is the parameter space.
 Given $\bm{\theta}$, the solution of the system, $x(t,\bm{s};\bm{\theta})$, is a space-time field,
-(1) $(t,\bm{s})\in[0,T]\times\mathbb{S}\mapsto x(t,\bm{s};\bm{\theta})$
+$(t,\bm{s})\in[0,T]\times\mathbb{S}\mapsto x(t,\bm{s};\bm{\theta})$ (1)
 where $\mathbb{S}$ is the spatial domain and $T>0$.
-Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),x(t,\bm{s}_{2};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}\in\mathbb{R}^{n}$ (say, using the Finite Element method), one can obtain the so-called snapshot data matrix, $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$, where $n_{T}$ is the number of temporal points.
-The POD method finds the optimal $r$-dimensional ($r\ll n$) vector subspace $\mathbb{W}\subset\mathbb{R}^{n}$ and an orthogonal projection, $\mathcal{W}:\mathbb{R}^{n}\rightarrow\mathbb{W}$ such that the following distance is minimized[12]
-(2) $\text{min}_{\mathcal{W}}\sum_{i=1}^{n_{T}}||\bm{x}(t_{i};\bm{\theta})-\mathcal{W}(\bm{x}(t_{i};\bm{\theta}))||^{2}$
+Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),x(t,\bm{s}_{2};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}$, one obtains the snapshot data matrix, $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$ with $n_{T}$ being the number of temporal points.
+The POD method finds the optimal $r$-dimensional ($r\ll n$) vector subspace $\mathbb{W}\subset\mathbb{R}^{n}$ and an orthogonal projection, $\mathcal{W}:\mathbb{R}^{n}\rightarrow\mathbb{W}$, such that the following distance is minimized
 […]
-Hence, we see both the advantages and limitations of the POD approach above.
-The optimal POD basis, which is obtained from the snapshot data generated at the parameter setting $\bm{\theta}$, is optimal and only optimal for parameter $\bm{\theta}$.
-As shown in Figure 1, if data are already available from numerically solving the full-order governing equation repeatedly at a set of pre-determined parameter settings, it is possible to obtain the optimal basis for each parameter setting using the existing approach.
-However, for any new parameter in the parameter space, new basis modes (which is optimal or near optimal for the new parameter) is needed so that one can adapt the ROM according to parameter changes.
-The lack of robustness against parameter change limits the applicability of ROMs for design, diagnosis, control, uncertainty quantification and real-time operations (e.g., digital twins), where parameter changes are common[1, 24, 25, 29].
-In this paper, we formulate the problem of adapting POD basis as a supervised statistical learning problem, for which the goal is to learn a mapping between the parameter $\bm{\theta}$ and POD basis $\bm{\Phi}$ that spans a low-dimensional vector subspace.
-This subspace regression problem can be stated as follows:
+where $||\cdot||=\sqrt{\left\langle\cdot,\cdot\right\rangle}$ and $\left\langle\cdot,\cdot\right\rangle$ is the inner product of the Hilbert space.
+[…]
+Hence, for parametric dynamical systems, it is critical for the POD modes to be sufficiently representative so that data generated from various parameter settings are well represented.
+This is particularly important for design, diagnosis, control, uncertainty quantification and real-time operations (e.g., digital twins) where parameter changes are common (Amsallem et al.,, 2009; Peherstorfer and Willcox, 2015a, ; Peherstorfer and Willcox, 2015b, ; Sargsyan et al.,, 2015).
+One approach, known as the greedy basis generation, finds the reduced basis by an iterative procedure (Hesthaven et al.,, 2016).
+At each iteration, one new basis function is added using one solution computed at a new parameter setting.
+Throughout multiple iterations, the overall precision is gradually enhanced and the final basis is rich enough to represent data from the parameter space.
+Alternatively, the optimal basis can be adapted as parameters change so that the POD basis can capture the variation of a system’s behavior over its parameter space (Bui-Thanh et al.,, 2008; Vetrano et al.,, 2011).
+In this paper, we focus on this approach and formulate the problem of adapting the POD basis as a supervised statistical learning problem.
+The goal is to learn a mapping between the parameter $\bm{\theta}$ and POD basis $\bm{\Phi}$ that spans a low-dimensional vector subspace.
+Such a subspace regression problem is stated as follows:
 Given $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$ for $\bm{\theta}_{i}\in\mathbb{P}$, the problem is to learn a mapping $\mathcal{T}:\mathbb{P}\rightarrow\mathcal{ST}(r,n)$ from the parameter space $\mathbb{P}\subset\mathbb{R}^{d}$ to a compact Stiefel manifold of matrices with orthogonal columns, which enables the prediction of the POD basis matrix $\bm{\Phi}^{*}$ given a new parameter $\bm{\theta}^{*}$ without numerically solving the computationally expensive full-order governing equation at $\bm{\theta}^{*}$ (see Figure 1).
@@
 As shown in Figure 1, when (training) data are available by numerically solving the full-order governing equation repeatedly at a (usually small) number of parameter settings $\{\bm{\theta}_{i}\}_{i=1}^{k}$, it is possible to obtain the optimal basis $\{\bm{\Phi}_{i}\}_{i=1}^{k}$ at each parameter setting, separately, using the existing approach.
-Once the proposed pGP regression learns the mapping between parameters and POD bases, a new basis can be generated given the new parameters, thus adapting the ROM for parameter changes (see Sec.3).
-Figure 1: A supervised statistical learning problem (based on the proposed Projected Gaussian Process) that establishes the mapping between parameters and POD basis modes so that ROMs can be adapted for parameter changes.
+Once the proposed pGP regression learns the mapping between parameters and POD bases, a new basis can be predicted given the new parameters, thus adapting the ROM for parameters change (see Sec.3).
+Figure 1: A supervised statistical learning problem based on the proposed pGP that enables the prediction of POD basis modes as parameters change.
 1.2 Literature Review
-Various strategies have been proposed to adapt ROM for parameter changes.
+Various approaches have been investigated for tackling the model reduction problem of parametric dynamical systems.
+Some examples include the interpolatory model reduction and rational interpolation (Baur et al.,, 2011; Gosea et al.,, 2021), multi-moment matching (Freund,, 2003; Benner and Feng,, 2014), balanced truncation (Wittmuess et al.,, 2016), matrix interpolation (Panzer et al.,, 2010; Geuss et al.,, 2013), etc.
+For a comprehensive and complete review of projection-based model reduction for parametric dynamical systems, we refer the readers to (Benner et al.,, 2015, 2017).
+In this section, we primarily focus on major strategies for adapting the POD basis for parameters change.
 Figure 2 presents one possible way to group existing approaches into four categories.
-This figure is used to facilitate the explanations of different approaches as well as how the proposed approach is built upon prior work.
+This figure is only used to facilitate the explanations of different approaches as well as how the proposed approach is built upon the prior work.
 The boundaries that separate different approaches may change if we examine these methods from different perspectives.
-$\bullet$ Strategy I involves pre-computing a database, i.e., a library, of reduced-order bases[1, 15].
+$\bullet$ Strategy I involves pre-computing a database, i.e., a library, of reduced-order bases (Amsallem et al.,, 2009; Kapteyn et al.,, 2020).
 For applications where this approach is feasible, ROM can be updated in real time over the set of parameters included in the database.
-$\bullet$ Strategy II computes the first-order total derivatives of the POD modes with respective to a parameter[8], and a new POD basis, given a small change of the parameters from a base point, can be found through the first-order expansion in the parametric space.
+$\bullet$ Strategy II computes the first-order total derivatives of the POD modes with respective to a parameter (Borggaard and Pelletier,, 2008), and a new POD basis, given a small change of the parameters from a base point, can be found through the first-order expansion in the parametric space.
 $\bullet$ Strategy III involves different approaches that aim to interpolate or predict new POD bases from the observed parameter-POD basis pairs.
-In particular, the POD basis interpolation on Grassmann Manifolds have been extensively investigated[2, 30, 6, 23].
-This approach is based on an important result that a subspace spanned by a POD basis is a point of the Grassmann Manifold that consists of $r$-dimensional subspaces in $\mathbb{R}^{n}$, where $n$ and $r$ are respectively the original and reduced-order dimensions.
-Hence, a tangent space, attached to a given base point in the Grassmann Manifold, can be constructed, and the subspaces from the Grassmann Manifold can be projected to the tangent space through a logarithmic map.
-After this critical step, interpolation can be done on the tangent space (such as the Lagrangian interpolation[2], inverse distance weighting[23], or Gaussian Process[13]), and the interpolated vector on the tangent space is eventually mapped back to the Grassmann Manifold through an exponential map.
-Hence, this method depends on the choice of the base point upon which the tangent space is attached to, and naturally requires the injective radius of the exponential map to be sufficiently large so that the approach is stable[12].
-In contrast, the subspace angle interpolation approach directly computes the geometric distance between two subspaces by finding the principal angels between two POD bases[34].
+In particular, the POD basis interpolation on Grassmann manifolds have been extensively investigated (Amsallem and Farhat,, 2008; Son,, 2013; Benner et al.,, 2015; Mosquera et al.,, 2019).
+This approach is based on an important result that a subspace spanned by a POD basis is a point on the Grassmann manifold that consists of $r$-dimensional subspaces in $\mathbb{R}^{n}$, where $n$ and $r$ are respectively the original and reduced-order dimensions.
+Hence, a tangent space, attached to a given base point on the Grassmann manifold, can be constructed, and the subspaces from the Grassmann manifold can be projected to the tangent space through a logarithmic map.
+After this critical step, interpolation can be done on the tangent space (such as the Lagrangian interpolation (Amsallem and Farhat,, 2008), inverse distance weighting (Mosquera et al.,, 2019), or Gaussian Process (Giovanis and Shields,, 2020)), and the interpolated vector on the tangent space is eventually mapped back to the Grassmann manifold through an exponential map.
+Hence, this method depends on the choice of the base point upon which the tangent space is attached to, and naturally requires the injective radius of the exponential map to be sufficiently large so that the approach is stable (Friderikos et al.,, 2020).
+In contrast, the subspace angle interpolation approach directly computes the geometric distance between two subspaces by finding the principal angels between two POD bases (Ye and Lim,, 2016).
 The principal angels can be viewed as a series of angles required to rotate one subspace to form another, and such a distance metric is used to interpolate subspaces.
+Other examples based on the idea of interpolation include the matrices interpolation of local reduced models for linear time-invariant systems (Panzer et al.,, 2010), and Son, (2013) investigated the interpolation on Grassmann manifolds for affinely dependent system matrices on model parameters.
 Figure 2: Existing strategies to adapt ROM for parameter changes
-$\bullet$ Strategy IV involves statistical machine learning to learn the mapping between parameters and POD bases (or the subspaces spanned by POD bases).
-Leveraging the key ideas from both the POD basis interpolation on Grassmann Manifold[2] and subspace angle interpolation[34], a regression tree-based approach on Grassmann Manifold was proposed in [18].
+$\bullet$ Strategy IV involves using statistical machine learning to learn the mapping between parameters and POD bases (or the subspaces spanned by POD bases).
+Leveraging the key ideas from both the POD basis interpolation on Grassmann manifold (Amsallem and Farhat,, 2008) and subspace angle interpolation (Ye and Lim,, 2016), a regression tree-based approach on Grassmann manifold was proposed in Liu and Liu, (2023).
 Following this approach, a tree is grown by repeatedly splitting the tree node (i.e., a binary partition the parameter space) to maximize the Riemannian distance between two subspaces (based on the principals angles) spanned by the predicted POD bases on the left and right daughter nodes.
-As a result, this tree-based approach effectively divides the parameter space into sub-regions, and find the optimal POD basis for each sub-region (which could generate useful insights on how a system’s behavior varies over the parameter space).
+As a result, this tree-based approach effectively divides the parameter space into sub-regions, and the POD basis for a sub-region is obtained using only the data generated from that sub-region.
 This approach successfully combines the POD basis prediction problem into the classical framework of regression trees, and appears to be more stable and conservative than the existing POD basis interpolation on Grassmann manifolds (as the tree-based method only updates the POD basis when the parameter jumps from one sub-region to another in the parameter space).
@@
 For those less important parameters, the tree-based approach effectively exclude those parameters from being used to split the tree (note that, the existing interpolation method does not automatically do this).
-Other statistical learning approaches involve the use of Gaussian Process (GP).
-[13] directly used GP to capture the relationship between parameters and the matrices that determine the vectors on the tangent space of the Grassmann Manifold.
+Finally, in very recent years, neural networks have also been used for adapting ROM.
+For example, Franco et al., (2024) proposed the Deep Orthogonal Decomposition by constructing a deep neural network to approximate the solution manifold through a continuously adaptive local basis (i.e., use a neural network to learn the mapping from parameters and subspaces).
+Berman and Peherstorfer, (2024) proposed to continuously adapt the low-rank weights of pre-train neural networks to predict the evolution of solution fields at new physics parameters.
+In general, training a neural network requires a large amount of training data (e.g., Franco et al., (2024) used 500 training samples for model reduction of the Eikonal equation).
+This may potentially pose additional challenges when it is expensive to generate training data by solving full-order physics.
+Other statistical learning approaches involve the use of GP.
+For example, Giovanis and Shields, (2020) directly used GP to capture the relationship between parameters and the matrices that determine the vectors on the tangent space of the Grassmann manifold.
 Hence, given a new parameter setting, GP is used to predict a new vector in the tangent space, which yields the predicted subspace on the Grassmann manifold through an exponential map.
 Note that, because GP performs prediction through a weighted linear combination of responses, this approach is similar to the existing Lagrangian POD basis interpolation on Grassmann manifolds but essentially with different interpolation weights.
-For the reasons noted in Sec.3, we refer this approach as the native GP approach following the terminology in [21].
-[20] proposed an approach known as Common POD (cPOD).
-In the context of engine injector design, a common POD basis is found for some base design geometry, the POD bases for other geometries are obtained by rescaling the common POD through a linear map.
+For the reasons noted in Sec.3, we refer to this approach as the native GP approach following the terminology in Mallasto and Feragen, (2018).
+Mak et al., (2018) proposed another approach known as the Common POD (cPOD).
+In an engine injector design example, a common POD basis is found for some base design geometry, the POD bases for other geometries are obtained by rescaling the common POD through a linear map.
 This approach requires sound domain knowledge of the system under consideration to determine the base design and establish the linear map between POD bases from the common POD.
-[35] proposed a GP subspace regression where the central idea is to directly construct a matrix-variate GP for the matrix bases such that the approach does not hinge on the differential geometric structures of the Gassmann manifold.
+Zhang et al., (2022) proposed a GP subspace regression where the central idea is to directly construct a matrix-variate GP for the matrix bases such that the approach does not hinge on the differential geometric structures of the Gassmann manifold.
 1.3 Contributions of this Paper
-The prior work in the literature has shown the great potential of leveraging machine learning for adapting projection-based ROMs.
-The main contribution of this paper is to propose a new supervised statistical learning method, known as the Projected Gaussian Process (pGP) regression, that enables the learning of the mapping between parameters and POD bases.
-To our best knowledge, this method is not yet available in the literature.
-Computer code is also available as Supplementary Materials.
-As a statistical machine learning approach, the proposed pGP provides some critical capabilities much needed for adapting ROMs against parameter changes:
+The prior work in the literature has shown the great potential of leveraging machine learning for adapting projection-based ROM.
+The main contribution of this paper is to propose a new supervised statistical learning method that enables the learning of the mapping between parameters and POD bases.
+To our best knowledge, this pGP approach is not yet available in the literature.
+As a statistical machine learning approach, the proposed pGP provides some critical capabilities much needed for adapting ROM against parameter changes:
 i).
-Unlike existing interpolation method (e.g., Lagrangian interpolation), the statistical learning approach allows us to optimally estimate (or tune) model parameters based on the training data (i.e., the prediction becomes problem-specific).
+Unlike existing interpolation methods (e.g., Lagrangian interpolation), the proposed statistical learning approach allows us to optimally estimate (or tune) model parameters based on training data (i.e., the predictions are data-driven).
 In contrast, when the Lagrangian interpolation is used, the Lagrange bases are computed from pre-existing equations which are independent of the problem or data of interest.
-Because the pGP model is always trained and optimized given the training dataset, such a flexibility potentially provides the opportunity to further improve the accuracy of ROM for the specific problem at hand (as shown in the numerical examples)
+Because the pGP model is always trained and optimized given a training dataset, such a flexibility provides the opportunity to further improve the accuracy of the predicted POD basis modes for specific problems at hand (as shown in the numerical examples)
 ii).
+The proposed pGP not only provides the point prediction of the POD basis, but also quantifies the statistical prediction uncertainty associated with the predicted POD modes.
+Such a capability of uncertainty quantification facilitates decision-making under uncertainty, and provides a critical link (as a future research path) to a large body of literature on statistical experimental design and deep reinforcement learning for more efficient construction of the POD basis by sequentially exploring the solution space of full-order models (Ivanova et al.,, 2021; Blau et al.,, 2022).
+iii).
 The proposed pGP, as a statistical learning approach, provides the capability of parameter selection when constructing the mapping between parameters and POD bases.
-This capability is largely missing in the literature but is particularly important for systems with high-dimensional parameter space, i.e., not all parameters are equally important in terms of how they change the POD basis modes.
-If relevant parameters can be identified, one only needs to explore a reduced parameter space, with a lower dimension, when investigating how system dynamics varies in the parameter space.
+This capability is needed when not all parameters are equally important in terms of their predictive power for POD modes.
+If important parameters can be identified, the statistical pGP model may require a smaller number of input.
+When experimental design is used to determine the optimal parameter settings for subsequent experiment, this can also potentially lower the dimension of the parameter space (i.e., the decision space in experimental design becomes smaller).
+We would also like to point out that this argument shares some similarities with the concept of active subspace (Constantine et al.,, 2014).
+The active subspace is usually a low-dimensional space of physics model input which contains the detected directions of the strongest variability in model output, while the parameter selection mentioned in this paper refers to the process of identifying the parameters which have stronger predictive power of the POD modes in a statistical model.
+Hence, the outcome of identifying the active subspace allows one to construct an approximation of the physics model on a low-dimensional subspace, while the outcome of parameter selection in this paper allows one to predict the POD modes using a smaller set of parameters.
 iii).
-Due to the parameter estimation/selection capabilities, the pGP approach naturally involves the use of global POD basis as its special case.
-When no parameter is identified as significant, the predicted POD basis matrices for different parameters, under the proposed pGP framework, become i.i.d.
-samples from a matrix-variate Gaussian distribution with the global POD basis matrix being the mean of this distribution (as shown in numerical Example II).
-iv).
-The proposed pGP not only provides the point prediction of the POD basis, but also scientifically quantifies the uncertainty associated with the predicted POD basis as well as the output generated by ROMs.
-Such a capability of uncertainty quantification facilitates decision-making under uncertainty, and provides a critical link (as a future research path) from the proposed approach to a large body of literature on statistical experimental design and deep reinforcement learning for more efficient construction of ROMs by sequentially exploring the solution space of full-order models[14, 7].
+Due to the parameter estimation/selection capabilities above, the pGP approach naturally involves the use of global POD basis as its special case.
+As shown in this paper, it is possible to set the parameters in the proposed pGP regression model such that the predicted POD basis matrices for different parameters become i.i.d.
+samples from a matrix-variate Gaussian distribution with the global POD basis matrix being the mean of this distribution (as shown in the numerical examples).
 The paper is organized as follows.
@@
 The proposed pGP approach is presented in Section 3.
-Comprehensive numerical examples and discussions are presented in Section 4.
+Numerical examples and discussions are presented in Section 4.
 Section 5 concludes the paper.
 2 Preliminaries on POD basis interpolation
-This section provides preliminaries on the interpolation problem of POD basis modes.
-As discussed in Section 1.1, due to the computational cost associated with solving full-order governing equations, it is only possible to obtain the snapshot matrices for a relatively small number of parameters, $\bm{\Lambda}=\{\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{k}\}$, and obtain the optimal POD basis $\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}$ for each parameter in $\bm{\Lambda}$.
-Hence, for any new parameter $\bm{\theta}^{*}\notin\bm{\Lambda}$, the interpolation of POD basis is needed to obtain $\bm{\Phi}^{*}$.
-Grassmann Manifold.
-In honour of Hermann Grassmann, a Grassmann Manifold, $\mathcal{G}(r,n):=\{\mathbb{W}\subset\mathbb{R}^{n},\mathrm{dim}(\mathbb{W})=r\}$, is a differentiable manifold that collects all $r$-dimensional linear subspaces in $\mathbb{R}^{n}$[5].
-Hence, the subspace $\mathbb{W}$ is a point in $\mathcal{G}(r,n)$ and the optimization problem in (2) essentially seeks the optimal point in $\mathcal{G}(r,n)$ that minimizes the distance $\sum_{i=1}^{n_{T}}||\bm{x}(t_{i};\bm{\theta})-\mathcal{W}(\bm{x}(t_{i};\bm{\theta}))||^{2}$.
+This section provides preliminaries on the interpolation of POD bases.
+Suppose that the snapshot data matrices are generated at a set of parameter settings, $\bm{\Lambda}=\{\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{k}\}$, and the optimal POD bases $\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}$ are obtained for each parameter in $\bm{\Lambda}$.
+Then, the POD basis interpolation problem involves finding the POD basis $\bm{\Phi}^{*}$ at new parameters $\bm{\theta}^{*}\notin\bm{\Lambda}$.
+Some important concepts are firstly introduced.
+[…]
 Stiefel manifold and a submersion.
 A subspace in $\mathcal{G}(r,n)$ is spanned by an orthogonal basis $\bm{\Phi}\in\mathbb{R}^{n\times r}$ ($\bm{\Phi}^{T}\bm{\Phi}=\bm{I}_{r}$).
-A collection of $n\times r$ matrices with orthogonal columns forms a compact Stiefel manifold denoted by $\mathcal{ST}(r,n)$[12].
+A collection of $n\times r$ matrices with orthogonal columns forms a compact Stiefel manifold denoted by $\mathcal{ST}(r,n)$.
 Note that, if $\bm{\Phi}$ spans a subspace $p\in\mathcal{G}(r,n)$, so does $\bm{\Phi}\bm{J}$ for any $r\times r$ matrix $\bm{J}$ such that $\bm{J}^{T}\bm{J}=\bm{I}_{p}$.
-Therefore, there exists is a fiber bundle
-(3) $\pi:\bm{\Phi}\in\mathcal{ST}(r,n)\mapsto\pi(\bm{\Phi})=p\in\mathcal{G}(r,n)$
-which indicates that any point $p\in\mathcal{G}(r,n)$ can be represented by a point of $\pi^{-1}(p)$.
+Therefore, there exists a fiber bundle
+$\pi:\bm{\Phi}\in\mathcal{ST}(r,n)\mapsto\pi(\bm{\Phi})=p\in\mathcal{G}(r,n)$ (3)
+which indicates that any point $p\in\mathcal{G}(r,n)$ can be represented by a point of the fiber $\pi^{-1}(p)$ (Friderikos et al.,, 2020).
 Riemannian metric.
@@
 For any two velocity vectors, $v_{1},v_{2}\in\mathcal{T}_{p}\mathcal{G}(r,n)$, we have
-(4) $\left\langle v_{1},v_{2}\right\rangle:=\left\langle\bm{Z}_{1},\bm{Z}_{2}\right\rangle=\mathrm{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$
-where $\bm{Z}_{1}$ and $\bm{Z}_{2}$ are $n\times r$ matrices from the horizontal space of $\bm{\Phi}=\pi^{-1}(p)$ denoted by $\mathbb{H}_{\bm{\Phi}}:=\{\bm{Z}\in\mathbb{R}^{n\times r};\bm{Z}^{T}\bm{\Phi}=0\}$.
-In other words, given a base point $p$, the Grassmann Manifold inherits the Riemannian structure of the Stiefel manifold due to (3) and becomes a Riemannian manifold.
+$\left\langle v_{1},v_{2}\right\rangle:=\left\langle\bm{Z}_{1},\bm{Z}_{2}\right\rangle=\mathrm{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$ (4)
+where $\bm{Z}_{1}$ and $\bm{Z}_{2}$ are $n\times r$ matrices from the horizontal space of $\bm{\Phi}\in\pi^{-1}(p)$ denoted by $\mathbb{H}_{\bm{\Phi}}:=\{\bm{Z}\in\mathbb{R}^{n\times r};\bm{Z}^{T}\bm{\Phi}=0\}$.
+In other words, given a base point $p$, the Grassmann manifold inherits the Riemannian structure of the Stiefel manifold due to (3) and is a Riemannian manifold.
 Exponential and Logarithm maps. For any point $p\in\mathcal{G}(r,n)$ and the attached tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, the Exponential map is defined as
-(5) $\text{Exp}_{p}:v\in\mathcal{T}_{p}\mathcal{G}(r,n)\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n)$
-where $\bm{\Phi}=\pi^{-1}(p)$ and $\bm{Z}=\bm{U}_{\bm{Z}}\bm{\Sigma}_{\bm{Z}}\bm{V}_{\bm{Z}}$ is the SVD of $\bm{Z}$.
-For any point $p\in\mathcal{G}(r,n)$ and a open set $U_{\bm{p}}:=\{p^{\prime}\in\mathcal{G}(r,n);\pi(\bm{\Phi})=p,\pi(\bm{\Phi}^{\prime})=p^{\prime},\pi(\bm{\Phi})^{T}\pi(\bm{\Phi}^{\prime})\text{ is invertible)}$), the Logarithm map is defined as
-(6) $\text{Log}_{p}:p^{\prime}\in U_{\bm{p}}\mapsto\text{Log}_{p}(p^{\prime})\in\mathcal{T}_{p}\mathcal{G}(r,n).$
+$\text{Exp}_{p}:v\in\mathcal{T}_{p}\mathcal{G}(r,n)\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n)$ (5)
+where $\bm{\Phi}\in\pi^{-1}(p)$ and $\bm{Z}=\bm{U}_{\bm{Z}}\bm{\Sigma}_{\bm{Z}}\bm{V}_{\bm{Z}}$ is the SVD of $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$.
+Note that, for a given $\Phi$ such that $\pi(\Phi)=p$, the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$ is isomorphic to the horizontal space $\mathbb{H}_{\bm{\Phi}}$.
+For any $v\in\mathcal{T}_{p}\mathcal{G}(r,n)$, there is a unique $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$ known as the horizontal lift of $v$ (Friderikos et al.,, 2020).
+For any point $p\in\mathcal{G}(r,n)$ and a open set $U_{\bm{p}}:=\{p^{\prime}\in\mathcal{G}(r,n)\}$, the Logarithm map is defined as
+$\text{Log}_{p}:p^{\prime}\in U_{\bm{p}}\mapsto\text{Log}_{p}(p^{\prime})\in\mathcal{T}_{p}\mathcal{G}(r,n).$ (6)
 Existing POD interpolation method.
-Within $U_{\bm{p}}$, the Logarithm map $\text{Log}_{p}$ is a diffemorphism, and the one-one correspondence between $v\in\mathcal{T}_{p}\mathcal{G}$ and $p^{\prime}\in U_{\bm{p}}$ can be defined through the Exponential and Logarithm maps[12].
-Hence, [2] described a well-known Lagrangian interpolation method for POD basis.
+Within $U_{\bm{p}}$, the Logarithm map $\text{Log}_{p}$ is a diffemorphism, and the one-to-one correspondence between $v\in\mathcal{T}_{p}\mathcal{G}$ and $p^{\prime}\in U_{\bm{p}}$ can be defined through the Exponential and Logarithm maps (Friderikos et al.,, 2020).
+Hence, Amsallem and Farhat, (2008) described a popular Lagrangian interpolation method for POD bases.
 In a nutshell, given the training data $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, subspaces corresponding to $\bm{\theta}_{i}$ are mapped to vectors $v_{i}$ in the tangent space through the Logarithm map.
 Then, the interpolation of vectors can be done in the tangent space, and the interpolated vector $v^{*}$ is mapped back to the Grassmann manifold through the Exponential map.
-Such a POD basis interpolation algorithm is provided in Algorithm 1.
-Algorithm 1 POD basis interpolation on Grassmann Manifold
-  Input: Integers $r$ and $n$ ($2r\leq n$), $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, and a target parameter $\bm{\theta}^{*}$.
+Algorithm 2.1 provides an example of such an approach for the one-dimensional case, i.e., $\theta\in\mathbb{R}$.
+ Input: Integers $r$ and $n$ ($2r\leq n$), $\{\theta_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, and a target parameter $\theta^{*}$.
   • Choose a matrix $\bm{\Phi}_{0}\in\{\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}\}$ such that $\bm{\Phi}_{0}^{T}\bm{\Phi}_{i}$ is non-singular for all $i$.
   • For each $i=1,2,\cdots,k$, perform the thin SVD on $\bm{\Phi}_{i}(\bm{\Phi}_{0}^{T}\bm{\Phi}_{i})^{-1}-\bm{\Phi}_{0}=\bm{U}_{i}\bm{\Sigma}_{i}\bm{V}_{i}^{T}$, and obtain a matrix $\bm{Z}_{i}$
-    (7) $\bm{Z}_{i}=\bm{U}_{i}\text{arctan}(\bm{\Sigma}_{i})\bm{V}_{i}^{T}.$
+    $\bm{Z}_{i}=\bm{U}_{i}\text{arctan}(\bm{\Sigma}_{i})\bm{V}_{i}^{T}.$ (7)
   • Compute an interpolated matrix and a thin SVD
-    (8) $\bm{Z}^{*}=\sum_{i=1}^{k}\prod_{i\neq j}\frac{\bm{\theta}^{*}-\bm{\theta}_{j}}{\bm{\theta}_{i}-\bm{\theta}_{j}}\bm{Z}_{i}=\bm{U}^{*}\bm{\Sigma}^{*}\bm{V}^{*}$
+    $\bm{Z}^{*}=\sum_{i=1}^{k}\prod_{i\neq j}\frac{\theta^{*}-\theta_{j}}{\theta_{i}-\theta_{j}}\bm{Z}_{i}=\bm{U}^{*}\bm{\Sigma}^{*}\bm{V}^{*}.$ (8)
+    Note that, when the dimension of the parameter space $d$ is larger than one, the Lagrangian interpolation in (8) can be replaced by other multidimensional interpolation methods.
   • Return an instability message if the largest singular value of $\bm{Z}^{*}$ is greater than $\pi/2$. Otherwise, compute a $n\times r$ matrix
-    (9) $\bm{\Phi}^{*}=\bm{\Phi}_{0}\bm{V}^{*}\text{cos}\bm{\Sigma}^{*}+\bm{U}^{*}\text{sin}\bm{\Sigma}^{*}$
+    $\bm{\Phi}^{*}=\bm{\Phi}_{0}\bm{V}^{*}\text{cos}\bm{\Sigma}^{*}+\bm{U}^{*}\text{sin}\bm{\Sigma}^{*}$ (9)
     where sin and cos only act on the diagonal of $\bm{\Sigma}^{*}$.
-  return  The interpolated POD basis $\bm{\Phi}^{*}$ for parameter $\bm{\theta}^{*}$.
-3 Projected Gaussian Process for Adapting ROMs
-In this section, we describe a supervised statistical learning method, based on the proposed pGP, for adapting ROMs against parameter changes.
-3.1 Projection from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$
+return  The interpolated POD basis $\bm{\Phi}^{*}$ for parameter $\theta^{*}$.
+Algorithm 1 POD basis interpolation on Grassmann manifold
+3 Projected Gaussian Process for Adapting ROM
+In this section, we present a supervised statistical learning method, based on the proposed pGP, for adapting POD bases against parameters change.
+3.1 A projection from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$
 We first show that there exists an injective mapping from the Euclidean space $\mathbb{R}^{nr-r}$ to the Grassmann manifold $\mathcal{G}(r,n)$.
@@
 However, it is important to point out that this seemingly natural approach needs to be carefully carried out because the matrix $\bm{Z}$ should be defined on the horizontal space $\mathbb{H}_{\bm{\Phi}}$ given a base point $p$, rather than the Euclidean space $\mathbb{R}^{n\times r}$.
-Precisely speaking, if $\bm{Z}$ is modeled as a GP with respect to the parameter $\bm{\theta}$ on $\mathbb{R}^{n\times r}$, then, the point prediction, $\hat{\bm{Z}}^{*}$, is given by a linear combination of $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ based on the well-known results of GP regression[28].
+Precisely speaking, if $\bm{Z}$ is modeled as a GP with respect to the parameter $\bm{\theta}$ on $\mathbb{R}^{n\times r}$, then, the point prediction, $\hat{\bm{Z}}^{*}$, is given by a linear combination of $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ based on the well-known results of GP regression (Rasmussen and Williams,, 2005).
 Since $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ are in the horizontal space $\mathbb{H}_{\bm{\Phi}}$ , the point prediction $\hat{\bm{Z}}^{*}$ remains in the same space.
 However, the uncertainty quantification or predictive internal of $\hat{\bm{Z}}^{*}$ (or, the predicted $\hat{\bm{\Phi}}^{*}$) is problematic because a sample $\bm{Z}$ from a Gaussian distribution on $\mathbb{R}^{n\times r}$ does not necessarily belong to $\mathbb{H}_{\bm{\Phi}}$.
-To remedy this issue, we first present the following results:
+To remedy this issue, we leverage the following result:
 Proposition 3.1.
-Given a basepoint $p\in\mathcal{G}(r,n)$ and $\bm{\Phi}=[\bm{\phi}_{1},\cdots,\bm{\phi}_{r}]$ such that $\pi(\bm{\Phi})=p$, let $\bm{y}$ be a $(nr-r)\times 1$ vector, $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ be a $r\times nr$ matrix with orthogonal rows, and $\bm{F}=[\bm{F}_{1}|\bm{F}_{2}|\cdots|\bm{F}_{r}]$ be a $(nr-r)\times nr$ matrix where each block $\bm{F}_{i}$, $i=1,2,\cdots,r$, is a $(nr-r)\times n$ matrix that satisfies the following conditions: (i) for any block $i$ ($i=1,2,\cdots,r$), only row $(i-1)\times(n-1)+1$ to row $i\times(n-1)$ contain non-zero entries; (ii) all non-zero row vectors are unit vectors; and (iii) all non-zero row vectors are orthogonal to each other (note that conditions (i)-(iii) imply $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$).
-Then,
-(10) $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{T}\bm{y})$
-is a $n\times r$ matrix that belongs to the horizontal space $\mathbb{H}_{\bm{\Phi}}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$.
-Proof. Note that $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$, i.e.,
-(11) $\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=0$
-where $\text{vec}(\bm{Z})$ is the column-wise vectorization of $\bm{Z}$, and $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ is a $r\times nr$ block diagonal matrix with orthogonal rows.
-Then, by introducing a $(nr-r)\times nr$ matrix $\bm{F}$ with orthogonal rows such that $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$, we have
-(12) $\begin{pmatrix}\tilde{\bm{\Phi}}\\ \bm{F}\end{pmatrix}\text{vec}(\bm{Z})=\begin{pmatrix}\bm{0}\\ \bm{y}\end{pmatrix}$
-where $\bm{y}$ is a $(nr-r)$-dimensional column vector.
-Because $\bm{F}\text{vec}(\bm{Z})=\bm{y}$ is an underdetermined system, we may find $\text{vec}(\bm{Z})$ by $\bm{F}^{\dagger}\bm{y}$ where $\bm{F}^{\dagger}=\bm{F}^{T}(\bm{F}\bm{F}^{T})^{-1}$.
-From conditions (i)-(iii) above, $\bm{F}\bm{F}^{T}$ is an identity matrix, and $\bm{F}^{\dagger}\bm{F}^{T}$ as shown in Proposition 3.1.
-Here, the choice of $\bm{F}$ depends on $\bm{\Phi}$ (and the basepoint $\bm{p}$ in Proposition 3.1) as it must satisfy conditions (i)-(iii).
-Let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}_{1},\tilde{\bm{\Phi}}_{2},\cdots,\tilde{\bm{\Phi}}_{r}]$, and let $\tilde{\bm{F}}=[\tilde{\bm{F}}_{1},\tilde{\bm{F}}_{2},\cdots,\tilde{\bm{F}}_{r}]$ be a $(nr-r)\times nr$ proposal matrix where each block $\tilde{\bm{F}}_{i}$ has the structure as specified by condition (i) above.
-Then, applying the Gram-Schmidt (G-S) process can be applied to the row vectors of each block $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \tilde{\bm{F}}_{i}\end{smallmatrix}\bigr{)}$ (for $i=1,2,\cdots,r$) and normalizing each non-zero rows after the G-S process, we obtain $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \bm{F}_{i}\end{smallmatrix}\bigr{)}$ where $\bm{F}_{i}$ satisfies conditions (i)-(iii).
-Note that, the G-S process starts from the $i$-th row because the leading $(i-1)$ rows of $\tilde{\bm{\Phi}}_{i}$ are zeros.
+For a given basepoint $p\in\mathcal{G}(r,n)$ and its corresponding basis matrix $\bm{\Phi}=[\bm{\phi}_{1},\cdots,\bm{\phi}_{r}]$, i.e, $\pi(\bm{\Phi})=p$, it is possible to obtain a $n\times r$ matrix $\bm{Z}$ from a $(nr-r)\times nr$ matrix $\bm{F}$ and a $(nr-r)\times 1$ vector $\bm{y}$ as follows
+$\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{T}\bm{y})$ (10)
+such that $\bm{Z}$ belongs to the horizontal space $\mathbb{H}_{\bm{\Phi}}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$.
+Here, the matrix $\bm{F}$ has orthogonal rows and satisfies the conditions $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$ and $\bm{F}\bm{F}^{T}=\bm{I}$, where $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ is a $r\times nr$ matrix with orthogonal rows.
+To prove Proposition 3.1, note that,
+$\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=0$ (11)
+when $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$.
+Then, if we can choose a $(nr-r)\times nr$ matrix $\bm{F}$ which has orthogonal rows and satisfies $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$, we have
+$\begin{pmatrix}\tilde{\bm{\Phi}}\\ \bm{F}\end{pmatrix}\text{vec}(\bm{Z})=\begin{pmatrix}\bm{0}\\ \bm{y}\end{pmatrix}$ (12)
+where $\bm{y}$ is a (nrr)(nr-r)-dimensional column vector.
+Because $\bm{F}\text{vec}(\bm{Z})=\bm{y}$ is an underdetermined system, one can find $\text{vec}(\bm{Z})$ by $\bm{F}^{\dagger}\bm{y}$ where $\bm{F}^{\dagger}=\bm{F}^{T}(\bm{F}\bm{F}^{T})^{-1}$.
+Next, if we further impose the condition $\bm{F}\bm{F}^{T}=\bm{I}$, then, $\bm{F}^{\dagger}=\bm{F}^{T}$ and $\text{vec}(\bm{Z})=\bm{F}^{T}\bm{y}$ as shown in Proposition 3.1.
+It is straightforward to verify that, when $\text{vec}(\bm{Z})=\bm{F}^{T}\bm{y}$, we have $\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=\tilde{\bm{\Phi}}\bm{F}^{T}\bm{y}=\bm{0}$, which implies that $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$ according to (11).
 Proposition 3.1 implies that, from a vector $\bm{y}\in\mathbb{R}^{nr-r}$, it is possible to compute a matrix $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ that comes from the horizontal space $\mathbb{H}_{\bm{\Phi}}$.
-In fact, it is not hard to understand why the dimension of $\bm{y}$ is $nr-r$.
-Although there are $n\times r$ elements in the matrix $\bm{Z}$, the constraint $\bm{Z}^{T}\bm{\Phi}=\bm{0}$ (i.e., $\bm{Z}$ must come from the horizontal space of $\bm{\Phi}$) takes $r$ degree of freedom away, leaving behind $nr-r$ degrees of freedom.
-Hence, for any basepoint $p\in\mathcal{G}(r,n)$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, we can define the following mapping from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$,
-(13) $\mathcal{P}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n).$
-Here, $\mathcal{P}_{p}:=\tilde{\mathcal{P}}_{p}\circ\text{Exp}_{p}$ where $\tilde{\mathcal{P}}_{p}$ maps an $(nr-r)$-dimensional vector to the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$,
-(14) $\tilde{\mathcal{P}}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto v\in\mathcal{T}_{p}\mathcal{G}(r,n)$
+In fact, it is not hard to understand why the dimension of $\bm{y}$ is nrrnr-r.
+Although there are $n\times r$ elements in the matrix $\bm{Z}$, the constraint $\bm{Z}^{T}\bm{\Phi}=\bm{0}$ consists of a number of $r$ linear equations (i.e., $\bm{Z}$ must come from the horizontal space of $\bm{\Phi}$) takes away $r$ degree of freedom, leaving behind nrrnr-r degrees of freedom.
+It is also noted that, the choice of $\bm{F}$ is obviously not unique as long as the matrix has orthogonal columns and satisfies $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$ and $\bm{F}\bm{F}^{T}=\bm{I}$.
+In our implementation, a computationally efficient procedure is used to construct the matrix $\bm{F}$.
+We let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}_{1},\tilde{\bm{\Phi}}_{2},\cdots,\tilde{\bm{\Phi}}_{r}]$, and let $\tilde{\bm{F}}=[\tilde{\bm{F}}_{1},\tilde{\bm{F}}_{2},\cdots,\tilde{\bm{F}}_{r}]$ be a $(nr-r)\times nr$ proposal matrix where each block $\tilde{\bm{F}}_{i}$ contains non-zero entries only from row $(i-1)\times(n-1)+1$ to row $i\times(n-1)$.
+Then, the Gram-Schmidt (G-S) process can be applied, in parallel, to the row vectors of each block $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \tilde{\bm{F}}_{i}\end{smallmatrix}\bigr{)}$ for $i=1,2,\cdots,r$.
+After the G-S process, each non-zero row is normalized, yielding $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \bm{F}_{i}\end{smallmatrix}\bigr{)}$ and $\bm{F}=[\bm{F}_{1},\bm{F}_{2},\cdots,\bm{F}_{r}]$ that satisfies the conditions needed, i.e., orthogonal rows, $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$ and $\bm{F}\bm{F}^{T}=\bm{I}$.
+Following Proposition 3.1, for any basepoint $p\in\mathcal{G}(r,n)$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, we can define the following mapping from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$,
+$\mathcal{P}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n).$ (13)
+Here, $\mathcal{P}_{p}:=\tilde{\mathcal{P}}_{p}\circ\text{Exp}_{p}$ where $\tilde{\mathcal{P}}_{p}$ maps an (nrr)(nr-r)-dimensional vector to the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$,
+$\tilde{\mathcal{P}}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto v\in\mathcal{T}_{p}\mathcal{G}(r,n)$ (14)
 and the Exponential map, $\text{Exp}_{p}$, is defined in (5).
 3.2 Injectivity of the map, $\mathcal{P}_{p}$
-Although $\mathcal{P}_{p}$ in (13) defines a map from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$, it is important to guarantee that such a map is injective (i.e., a subspace in $\mathcal{G}(r,n)$ can be uniquely determined by a vector in $\Omega_{p}$).
+Although $\mathcal{P}_{p}$ in (13) defines a map from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$, it is important to guarantee that such a map is injective.
 Proposition 3.2 below gives the condition for which $\mathcal{P}_{p}$ is injective.
-Proposition 3.2.
-Given a basepoint $p\in\mathcal{G}(r,n)$, there exists an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$
-(15) $\Omega_{p}:=\left\{\bm{y}\in\mathbb{R}^{nr-r};||\bm{y}||<\frac{\pi}{2}\right\}$
+Proposition 3.2. Given a basepoint $p\in\mathcal{G}(r,n)$, there exists an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$
+[…]
 such that the mapping $\mathcal{P}_{p}$ is injective.
-Proof. Consider an open set $D_{p}\subset\mathcal{T}_{p}\mathcal{G}(r,n)$
-(16) $D_{p}:=\left\{v\in\mathcal{T}_{p}\mathcal{G}(r,n);||v||<\pi/2\right\}$
-where $\pi/2$ is the injective radius on the Grassmann Manifold[12].
+To prove Proposition 3.2, consider an open set $D_{p}\subset\mathcal{T}_{p}\mathcal{G}(r,n)$
+[…]
+where $\pi/2$ is the injective radius on the Grassmann manifold (Friderikos et al.,, 2020).
 Hence, we seek an open set $\Omega_{p}\subset\mathbb{R}^{nr-r}$ such that $\mathcal{A}_{p}$ maps a vector $\bm{y}\in\Omega_{p}$ to the open disk $D_{p}$ above.
 Because $\bm{Z}=\text{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$, we have
-(17) $\begin{split}\bm{Z}^{T}\bm{Z}&=\begin{bmatrix}\bm{y}^{T}\bm{F}_{1}\\ \bm{y}^{T}\bm{F}_{2}\\ \vdots\\ \bm{y}^{T}\bm{F}_{r}\end{bmatrix}\begin{bmatrix}\bm{F}_{1}^{T}\bm{y}&\bm{F}_{2}^{T}\bm{y}&\cdots&\bm{F}_{r}^{T}\bm{y}\end{bmatrix}\\ &=\begin{bmatrix}\sum_{i=1}^{n-1}y_{(i-1)\times r+1}^{2}&&&\\ &\sum_{i=1}^{n-1}y_{(i-1)\times r+2}^{2}&&\\ &&\ddots&\\ &&&\sum_{i=1}^{n-1}y_{(i-1)\times r+r}^{2}\end{bmatrix}\end{split}$
+$\begin{split}\bm{Z}^{T}\bm{Z}&=\begin{bmatrix}\bm{y}^{T}\bm{F}_{1}\\ \bm{y}^{T}\bm{F}_{2}\\ \vdots\\ \bm{y}^{T}\bm{F}_{r}\end{bmatrix}\begin{bmatrix}\bm{F}_{1}^{T}\bm{y}&\bm{F}_{2}^{T}\bm{y}&\cdots&\bm{F}_{r}^{T}\bm{y}\end{bmatrix}\\ &=\begin{bmatrix}\sum_{i=1}^{n-1}y_{(i-1)\times r+1}^{2}&&&\\ &\sum_{i=1}^{n-1}y_{(i-1)\times r+2}^{2}&&\\ &&\ddots&\\ &&&\sum_{i=1}^{n-1}y_{(i-1)\times r+r}^{2}\end{bmatrix}\end{split}$ (17)
 Because the Riemannian metric on $\mathcal{G}(r,n)$ is defined by $\left\langle v_{1},v_{2}\right\rangle_{p}=\text{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$ given the basepoint $p$, we have
-(18) $||v||=\sqrt{\text{trace}(\bm{Z}^{T}\bm{Z})}=\left(\sum_{i=1}^{nr-r}y_{i}^{2}\right)^{\frac{1}{2}}=||\bm{y}||$
-which needs to be smaller than $\pi/2$. Hence, we have an open set $\Omega_{p}$, defined in (15), such that $\mathcal{P}_{p}$ is injective.
-3.3 Prediction based on the pGP regression
-Proposition 3.1 suggests an effective approach to establish the mapping between the parameter and POD basis.
+[…]
+which needs to be smaller than $\pi/2$ according to (16). Hence, we have an open set $\Omega_{p}$, defined in (15), such that $\mathcal{P}_{p}$ is injective.
+3.3 Prediction using the pGP regression
+Propositions 3.1 and 3.2 suggest an effective approach to establish the mapping between the parameter and POD basis.
 In a nutshell, as illustrated in Figure 3, a GP can be established to describe the mapping between the parameters space $\mathbb{R}^{d}$ and $\mathbb{R}^{nr-r}$, and then, the GP is projected to the Grassmann manifold.
 As a result, for any new parameter $\bm{\theta}^{*}$, a vector $\bm{y}^{*}$ can be found and mapped back to the Grassmann manifold to determine the optimal subspace onto which the full-order system is projected (equivalently, the POD basis matrix $\bm{\Phi}^{*}$).
-Figure 3: A sketch of the key relationships between different spaces/manifolds involved in the pGP regression.
-Let the vector $\bm{y}$ follows the multivariate Gaussian distribution, i.e., $\bm{y}\sim\mathcal{N}(\bm{0}_{nr-r,1},\bm{K})$ where $\bm{K}$ is a $(nr-r)\times(nr-r)$ symmetric and positive semi-definite matrix.
+Let the vector $\bm{y}$ follow a multivariate Gaussian distribution, i.e., $\bm{y}\sim\mathcal{N}(\bm{0}_{nr-r,1},\bm{K})$ where $\bm{K}$ is a $(nr-r)\times(nr-r)$ symmetric and positive semi-definite matrix.
 Then, $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ follows a matrix-variate Gaussian, $\bm{Z}\sim\mathcal{N}(\bm{0}_{n\times r},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$ in the sense that $\text{vec}(\bm{Z})\sim\mathcal{N}(\bm{0}_{nr\times 1},\bm{F}^{\dagger}\bm{K}(\bm{F}^{\dagger})^{T})$.
 For a given point $\mu\in\mathcal{G}(r,n)$ and a symmetric positive semi-definite matrix $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, we say that a random point $X\in\mathcal{G}(r,n)$ follows a projected Gaussian Distribution (pGD) on $\mathcal{G}(r,n)$
-(19) $X\sim\mathcal{P}_{\mu}(\mathcal{N}(\bm{0}_{nr-r,1},\bm{K}))$
+$X\sim\mathcal{P}_{\mu}(\mathcal{N}(\bm{0}_{nr-r,1},\bm{K}))$ (19)
 where the projection, $\mathcal{P}_{\mu}:\mathbb{R}^{nr-r}\rightarrow\mathcal{G}(r,n)$, is associated with the given point $\mu$.
 We denote the pGD by $\mathcal{N}_{\mathcal{G}}(\mu,\bm{K})$, and write $X\sim\mathcal{N}_{\mathcal{G}}(\mu,\bm{K})$, $\mu:=\mu_{\mathcal{G}}(X)\in\mathcal{G}(r,n)$, and $\text{cov}_{\mathcal{G}}(X):=\bm{K}$.
+Figure 3: A sketch of the key relationships between different spaces/manifolds involved in the pGP regression.
 To obtain the mapping $\mathcal{P}_{\mu}$ that connects $\mathbb{R}^{nr-r}$ and $\mathcal{G}(r,n)$ given a point $\mu\in\mathcal{G}(r,n)$ and $\bm{\Phi}=\pi^{-1}(\mu)\in\mathcal{ST}(r,n)$, we first introduce an operator $\mathcal{P}_{1}$ such that $\mathcal{P}_{1}(\bm{y})=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})=\bm{Z}$.
@@
 Hence, the mapping $\mathcal{P}_{\mu}$ can be defined as
-(20) $\mathcal{P}_{\mu}:=\text{span}\circ\mathcal{P}_{2}\circ\mathcal{P}_{1}$
+$\mathcal{P}_{\mu}:=\text{span}\circ\mathcal{P}_{2}\circ\mathcal{P}_{1}$ (20)
 that maps a $(nr-r)\times 1$ vector from $\mathbb{R}^{nr-r}$ to a $r$-dimensional subspace in $\mathcal{G}(r,n)$.
-Proposition 3.3.
-If $X_{i}\sim\mathcal{N}_{\mathcal{G}_{i}}(\mu_{i},\bm{K}_{i})$ for $i=1,2$, then, $(X_{1},X_{2})$ are jointly PGD on $\mathcal{G}_{1}\times\mathcal{G}_{2}$
-(21) $(X_{1},X_{2})\sim\mathcal{N}_{\mathcal{G}_{1}\times\mathcal{G}_{2}}\left(\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix},\tilde{\bm{K}}\right),\quad\quad\tilde{\bm{K}}=\begin{pmatrix}\bm{K}_{1},\bm{K}_{12}\\ \bm{K}_{12}^{T},\bm{K}_{2}\end{pmatrix}$
+Proposition 3.3. If $X_{i}\sim\mathcal{N}_{\mathcal{G}_{i}}(\mu_{i},\bm{K}_{i})$ for $i=1,2$, then, $(X_{1},X_{2})$ are jointly pGD on $\mathcal{G}_{1}\times\mathcal{G}_{2}$
+$(X_{1},X_{2})\sim\mathcal{N}_{\mathcal{G}_{1}\times\mathcal{G}_{2}}\left(\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix},\tilde{\bm{K}}\right),\quad\quad\tilde{\bm{K}}=\begin{pmatrix}\bm{K}_{1},\bm{K}_{12}\\ \bm{K}_{12}^{T},\bm{K}_{2}\end{pmatrix}$ (21)
 Then, the conditional distribution of $X_{1}$ given $X_{2}=p_{2}$ is given by
-(22) $X_{1}\mid(X_{2}=p_{2})\sim\mathcal{P}_{\mu_{1}}(\mathcal{MN}(\bm{u}_{\bm{s}},\bm{K}_{\bm{s}}))$
+$X_{1}\mid(X_{2}=p_{2})\sim\mathcal{P}_{\mu_{1}}(\mathcal{MN}(\bm{u}_{\bm{s}},\bm{K}_{\bm{s}}))$ (22)
 where $\bm{u}_{\bm{s}}=\bm{K}_{12}\bm{K}_{2}^{-1}\bm{s}$, $\bm{K}_{\bm{s}}=\bm{K}_{1}-\bm{K}_{12}\bm{K}_{2}^{-1}\bm{K}_{12}^{T}$, $\bm{s}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$, and $\mathcal{P}_{\mu_{2}}^{-1}$ is the inverse map.
-Proposition 3.3 is obtained following a similar technique in [21].
+Proposition 3.3 is obtained following a similar technique in Mallasto and Feragen, (2018).
 Let $\mathbb{A}_{1}=\mathcal{P}_{\mu_{1}}^{-1}(p_{1})=\{\bm{y}\in\mathbb{R}^{nr-r}:\mathcal{P}_{\mu_{1}}(\bm{Y})=p_{1}\}$ be the preimage of $p_{1}$ in $\mathbb{R}^{nr-r}$ associated with $\bm{\Phi}_{\mu_{1}}$ given a point $\mu_{1}$, and let $\mathbb{A}_{2}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$ similarly be the preimage of $p_{2}$, then,
-(23) $\begin{split}\text{Pr}(X_{1}|X_{2}=p_{2})&=\frac{\text{Pr}(\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2})}{\text{Pr}(\bm{y}_{2}\in\mathbb{A}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\frac{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}{\text{Pr}(\mathbb{A}_{2})}\frac{\mathcal{MN}((\bm{y}_{1}^{T},\bm{y}_{2}^{T})^{T};\bm{0},\tilde{\bm{K}})}{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\lambda_{\bm{y}_{2}}\mathcal{MN}(\bm{y}_{1};\bm{u}_{\bm{y}_{2}},\bm{K}_{\bm{y}_{2}})\end{split}$
+$\begin{split}\text{Pr}(X_{1}|X_{2}=p_{2})&=\frac{\text{Pr}(\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2})}{\text{Pr}(\bm{y}_{2}\in\mathbb{A}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\frac{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}{\text{Pr}(\mathbb{A}_{2})}\frac{\mathcal{MN}((\bm{y}_{1}^{T},\bm{y}_{2}^{T})^{T};\bm{0},\tilde{\bm{K}})}{\mathcal{MN}(\bm{y}_{2};\bm{0},\bm{K}_{2})}\\ &=\sum_{\bm{y}_{1}\in\mathbb{A}_{1},\bm{y}_{2}\in\mathbb{A}_{2}}\lambda_{\bm{y}_{2}}\mathcal{MN}(\bm{y}_{1};\bm{u}_{\bm{y}_{2}},\bm{K}_{\bm{y}_{2}})\end{split}$ (23)
 where $\bm{u}_{\bm{y}_{2}}=\bm{K}_{12}\bm{K}_{2}^{-1}\bm{y}_{2}$, and $\bm{K}_{\bm{y}_{2}}=\bm{K}_{1}-\bm{K}_{12}\bm{K}_{2}^{-1}\bm{K}_{12}^{T}$.
 Hence,
-(24) $X_{1}\mid(X_{2}=p_{2})\sim\mathcal{P}_{\mu_{1}}(\mathcal{MN}(\bm{u}_{\bm{s}},\bm{K}_{\bm{s}}))$
+$X_{1}\mid(X_{2}=p_{2})\sim\mathcal{P}_{\mu_{1}}(\mathcal{MN}(\bm{u}_{\bm{s}},\bm{K}_{\bm{s}}))$ (24)
 where $\bm{s}=\mathcal{P}_{\mu_{2}}^{-1}(p_{2})$.
-Once a PGD is obtained as described above, a Projected Gaussian Process (pGP) on Grassmann Manifold can be defined.
+Once a pGD is obtained as described above, a pGP on Grassmann manifold can be defined.
 We start with a multivariate GP defined in the Euclidean space.
 A collection $\bm{y}$ of random vectors is a Multivariate GP (MGP) in $\mathbb{R}^{d}$, denoted by $\bm{y}\sim\mathcal{MGP}(f,\omega,\bm{K})$, with a matrix-valued mean function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{nr-r}$, kernel $\omega:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, if any finite collection of $\bm{y}_{1}$, $\bm{y}_{2}$, $\cdots$, $\bm{y}_{k}$ have a joint multivariate Gaussian distribution such that
-(25) $[\bm{y}_{1}^{T},\bm{y}_{2}^{T},\cdot,\bm{y}_{k}^{T}]^{T}\sim\mathcal{N}([f^{T}(\bm{\theta}_{1}),f^{T}(\bm{\theta}_{2}),\cdot,f^{T}(\bm{\theta}_{k})^{T}]^{T},\bm{\Omega}\otimes\bm{K})$
-where $\bm{\Omega}\in\mathbb{R}^{k\times k}$ with $\bm{\Omega}_{i,j}=\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$. Then, a pGP on $\mathcal{G}(r,n)$ can be defined as follows:
-Projected Gaussian Process. A Projected Gaussian Process (pGP) on $\mathcal{G}(r,n)$ is defined by
-(26) $g\sim\mathcal{PGP}(m,\omega,\bm{K})$
+$\bm{y}_{1}^{T},\bm{y}_{2}^{T},\cdot,\bm{y}_{k}^{T}]^{T}\sim\mathcal{N}([f^{T}(\bm{\theta}_{1}),f^{T}(\bm{\theta}_{2}),\cdot,f^{T}(\bm{\theta}_{k})^{T}]^{T},\bm{\Omega}\otimes\bm{K})$ (25)
+where $\bm{\Omega}\in\mathbb{R}^{k\times k}$ with $\bm{\Omega}_{i,j}=\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$. Then, a pGP on $\mathcal{G}(r,n)$ can be defined.
+Projected Gaussian Process. A pGP on $\mathcal{G}(r,n)$ is defined by
+$g\sim\mathcal{PGP}(m,\omega,\bm{K})$ (26)
 where a collection $(g(\bm{\theta}_{1}),g(\bm{\theta}_{2}),\cdots,g(\bm{\theta}_{k}))\sim\mathcal{P}_{m}(\mathcal{MGP}(\bm{0},\omega,\bm{K})$ (i.e., a joint pGD), $m(\bm{\theta}):=\mu_{\mathcal{G}}(g(\theta))$ is the basepoint function and $k(\bm{\theta}_{i},\bm{\theta}_{i}):=\text{cov}_{\mathcal{G}}(g(\bm{\theta}_{i}),g(\bm{\theta}_{j}))=\omega(\bm{\theta}_{i},\bm{\theta}_{j})\bm{K}$ is the co-variance function in $\mathbb{R}^{nr-r}$.
@@
 Following Proposition 3.3, the joint distribution between the training output $\bm{p}^{\text{train}}=(p_{1},\cdots,p_{k})^{T}$ and test output $p_{*}$ is given by:
-(27) $\begin{pmatrix}p_{*}\\ \bm{p}^{\text{train}}\end{pmatrix}\sim\mathcal{N}_{\mathcal{G}_{*}\times\mathcal{G}_{1}\times\cdots\times\mathcal{G}_{k}}\left(\begin{pmatrix}m_{*}\\ \bm{m}^{\text{train}}\end{pmatrix},\begin{pmatrix}\bm{K}_{**},\bm{K}_{*}\\ \bm{K}_{*}^{T},\bm{K}^{\text{train}}\end{pmatrix}\right)$
+$\begin{pmatrix}p_{*}\\ \bm{p}^{\text{train}}\end{pmatrix}\sim\mathcal{N}_{\mathcal{G}_{*}\times\mathcal{G}_{1}\times\cdots\times\mathcal{G}_{k}}\left(\begin{pmatrix}m_{*}\\ \bm{m}^{\text{train}}\end{pmatrix},\begin{pmatrix}\bm{K}_{**},\bm{K}_{*}\\ \bm{K}_{*}^{T},\bm{K}^{\text{train}}\end{pmatrix}\right)$ (27)
 where $m_{*}=m(\theta_{*})$, $\bm{m}^{\text{train}}=(m(\bm{\theta}_{1}),m(\bm{\theta}_{2}),\cdots,m(\bm{\theta}_{d}))^{T}$, $\bm{K}^{\text{train}}=\bm{\Omega}\otimes\bm{K}\in\mathbb{R}^{k(nr-r)\times k(nr-r)}$, $K_{**}=\omega(\bm{\theta}_{*},\bm{\theta}_{*})\bm{K}\in\mathbb{R}^{(nr-r)\times(nr-r)}$, $\bm{K}_{*}=(\omega(\bm{\theta}_{*},\bm{\theta}_{1}),\omega(\bm{\theta}_{*},\bm{\theta}_{2}),\cdots,\omega(\bm{\theta}_{*},\bm{\theta}_{d}))\otimes\bm{K}\in\mathbb{R}^{(nr-r)\times k(nr-r)}$.
 Finally,
-(28) $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}))$
+$)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}$)) (28)
 where
-(29) $\bm{u}_{*}=\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}((\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T},(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T},\cdots,(\mathcal{P}_{m(\bm{\theta}_{d})}^{-1}(p_{k}))^{T})^{T}$
+$\bm{u}_{*}=\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}((\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T},(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T},\cdots,(\mathcal{P}_{m(\bm{\theta}_{d})}^{-1}(p_{k}))^{T})^{T}$ (29)
 and
-(30) $\tilde{\bm{K}}_{*}=\bm{K}_{**}-\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}\bm{K}_{*}^{T}.$
-Hence, a point estimate of the subspace $p_{*}\in\mathcal{G}(r,n)$ can be obtained by $\hat{p}_{*}=\mathcal{P}_{m_{*}}(\bm{u}_{*})$.
-Because $\bm{u}_{*}$ is the mean of the multivariate Gaussian distribution $\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*})$ defined in $\mathcal{R}^{nr-r}$, $\hat{p}_{*}$ is a maximum a posteriori probability (MAP) estimate.
-Also recall that $||\bm{u}_{*}||$ needs to be smaller than $\pi/2$ for $\mathcal{P}_{m_{*}}$ to be injective.
-Hence, if $||\bm{u}_{*}||>\pi/2$, we shrink this vector $\bm{u}_{*}\leftarrow\frac{\pi}{2}\frac{\bm{u}_{*}}{||\bm{u}_{*}||}$, which is the projection of the original vector $\bm{u}_{*}$ onto the surface of a $(nr-r)$-dimensional sphere.
-Similarly, the estimate $\hat{\bm{\Phi}}_{*}$ is obtained as $\mathcal{P}_{2}\circ\mathcal{P}_{1}(\bm{u}_{*})$ such that $\hat{p}_{*}$ is spanned by $\hat{\bm{\Phi}}_{*}$.
+$\tilde{\bm{K}}_{*}=\bm{K}_{**}-\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}\bm{K}_{*}^{T}.$ (30)
+[…]
 Figure 4: An illustration of the subspace prediction using the pGP regression.
-Figure 4 illustrate the process of predicting the optimal subspace using pGP.
-In this illustration, suppose that we obtain the snapshot data from three parameter settings $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$, as well as the optimal subspaces $p_{1}$, $p_{2}$ and $p_{3}$ that can be found using POD.
-The prediction of optimal subspace involves the following steps:
+Figure 4 illustrates the process of predicting the optimal subspace using pGP.
+In this illustration, suppose that we obtain the snapshot data from three parameter settings $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$, as well as the optimal subspaces $p_{1}$, $p_{2}$ and $p_{3}$ spanned by the POD modes found at $\bm{\theta}_{1}$, $\bm{\theta}_{2}$ and $\bm{\theta}_{3}$.
+The prediction of the optimal subspace involves the following steps:
 (i) Given the mean function $m(\bm{\theta})$, the three subspaces $p_{1}$, $p_{2}$ and $p_{3}$ are firstly mapped to the Euclidean space $\mathbb{R}^{nr-r}$ through the projection $\mathcal{P}^{-1}_{m(\bm{\theta})}$; see (29).
@@
 (ii) The three column vectors $(\mathcal{P}_{m(\bm{\theta}_{1})}^{-1}(p_{1}))^{T}$, $(\mathcal{P}_{m(\bm{\theta}_{2})}^{-1}(p_{2}))^{T}$ and $(\mathcal{P}_{m(\bm{\theta}_{3})}^{-1}(p_{3}))^{T}$ are used to predict a new column vector $\bm{u}_{*}$, for the new parameter $\bm{\theta}_{*}$, through a linear combination defined by a GP; see (29).
-(iii) The covariance matrix $\tilde{\bm{K}}_{*}$ is computed on $\mathbb{R}^{nr-r}$, which is exactly the same as the convectional GP.
-Steps (ii) and (iii) together yield a multivariate Gaussian distribution of $\bm{y}_{*}$) on $\mathbb{R}^{nr-r}$, i.e., $\bm{y}_{*}\sim\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*})$.
-(iv) Finally, the Gaussian distribution of $\bm{y}_{*}$ on $\mathbb{R}^{nr-r}$ (or, the predicted column vector $\bm{u}_{*}$ if only a point prediction is needed), is projected to the Grassmann Manifold through the projector $\mathcal{P}_{m(\bm{\theta}_{*})}$.
+[…]
+(iv) Finally, the Gaussian distribution of $\bm{y}_{*}$ on $\mathbb{R}^{nr-r}$ (or, the predicted column vector $\bm{u}_{*}$ if only a point prediction is needed) is projected back to the Grassmann manifold through the projector $\mathcal{P}_{m(\bm{\theta}_{*})}$.
 3.4 Discussions on model specification and parameter estimation
-A close examination of (28) reveals an important difference between how prediction is done by conventional GP and the proposed pGP.
-Using a conventional GP, the point predictor is a linear combination of the observations and the weights are determined by the covariance function (i.e., the point prediction does not depend the mean of the GP; a property well-known in the literature).
+A close examination of (28) reveals an important difference between how predictions are made by conventional GP and the proposed pGP.
+Using a conventional GP, the point predictor is a linear combination of the observations and the weights are determined by the covariance function (i.e., the point prediction does not depend on the mean of the GP; a property well-known in the literature).
 For the proposed pGP, as indicated by (28), the point prediction does depend on the function $m(\bm{\theta})$ that describes the relationship between the parameter and basepoint on the Grassmann manifold.
@@
 which can be obtained using a global POD basis.
-The second approach is to use another regression model to determine $m(\bm{\theta})$, such as the tree-based method described in [18].
+The second approach is to use another regression model to determine $m(\bm{\theta})$, such as the tree-based method described in Liu and Liu, (2023).
 In addition, because $n$ is usually large, a practical way to parameterize the $(nr-r)\times(nr-r)$ covariance matrix $\bm{K}$ is to let $\bm{K}=\sigma_{\bm{K}}\bm{I}$ and estimate $\sigma_{\bm{K}}$ from data.
 As for the kernel function $\omega(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\xi})$, existing Gaussian, Exponential or other kernel functions can be used, and the parameter $\bm{\xi}$ can be estimated from data.
-Hence, once $m(\bm{\theta})$ is determined, the cross-validation-based approach can be used to estimate $(\sigma_{\bm{K}},\bm{\xi})$ by minimizing $\sum_{i=1}^{k}\delta_{i}$, where the cross validation error is given by $\delta_{i}=||\bm{D}(\bm{\theta}_{i})-\hat{\bm{\Phi}}_{-i}\hat{\bm{\Phi}}_{-i}^{T}\bm{D}(\bm{\theta}_{i})||_{F}^{2}$, and $\hat{\bm{\Phi}}_{-i}$ is the estimated $\bm{\Phi}_{i}$ excluding data $\bm{D}(\bm{\theta}_{i})$.
-Also note that, for systems with high-dimensional parameter space (i.e., relatively large $d$), it is often the case that not all parameters are equally important in terms of how they affect the POD bases.
-In other words, the ROM can be sensitive to some parameters while independent of the others.
-Hence, when constructing the statistical learning model that establishes the mapping from parameter space to Grassmann Manifold, it is necessary to identify those relevant parameters.
-For the pGP model, one can achieve this goal using the Automatic Relevance Determination (ARD)[28].
+Hence, once $m(\bm{\theta})$ is determined, $(\sigma_{\bm{K}},\bm{\xi})$ can be estimated using existing methods such as the cross-validation-based approach or maximum likelihood estimation described in Rasmussen and Williams, (2005).
+Also note that, for systems with high-dimensional parameter spaces (i.e., relatively large $d$), it is often the case that not all parameters are equally important in terms of their predictive power of the POD modes.
+In other words, the POD modes can be sensitive to some parameters while independent of the others.
+Hence, when constructing the statistical learning model that establishes the mapping from parameter space to Grassmann manifold, it is useful to identify those relevant parameters.
 To illustrate the idea, consider the squared Exponential kernel given by
-(31) $\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}$
+$\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}$ (31)
 where $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{4}^{-2},\cdots,\xi_{d+2}^{-2}\}$.
-Here, the parameters $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ play a role of characteristic length-scales for parameter $\theta_{1},\theta_{2},\cdots,\theta_{d}$.
+Here, the parameters $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ play the role of characteristic length-scales for parameter $\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{d}$.
 If the estimated length-scales are very large, the corresponding parameters have little effects on the prediction and are practically removed from the model.
+This is known as the Automatic Relevance Determination (ARD) (Rasmussen and Williams,, 2005) and more details can be found in Sec.4.
 Finally, it is worth noting that the proposed pGP naturally involves the use of the global POD basis as its special case.
+Take the squared Exponential kernel above as an example.
 If $\xi_{1}=0$ and the estimated $\xi_{3},\xi_{4},\cdots,\xi_{d+2}$ are all close to zero, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ becomes practically zero.
-In this case, (29) immediately implies that $\bm{u}_{*}=\bm{0}$ and $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}))$.
-If the global POD basis is chosen as the reference point (which spans a subspace in the Grassmann Manifold where the tangent space is attached to) and the basepoint function $\bm{m}$ is invariant with $\bm{\theta}$, then, any subspace, $p_{i}$, corresponding to the ROM at parameter $\bm{\theta}_{i}$ on the Grassmann Manifold can be seen as i.i.d.
-samples from a PGD, $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}))$.
+In this case, (29) immediately implies that $\bm{u}_{*}=\bm{0}$ and $)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}$)).
+If the global POD basis is chosen as the reference point (which spans a subspace in the Grassmann manifold where the tangent space is attached to) and the basepoint function $\bm{m}$ is invariant with $\bm{\theta}$, then, any subspace $p_{i}$, corresponding to the parameter $\bm{\theta}_{i}$, on the Grassmann manifold can be seen as i.i.d.
+samples from a pGD, $)\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}$)).
 Hence, the point prediction, $\hat{p}_{*}$, is just the subspace spanned by the global POD basis.
+More details can be found in Sec.4.
 4 Numerical Examples
-Three numerical experiments are performed to demonstrate the proposed method: (i) flow around a circular cylinder; (ii) heat transfer in powder bed fusion additive manufacturing; and (iii) advection-diffusion processes.
-4.1 Example I: Flow around a Cylinder
-We first introduce the experiment setup and then present the experiment results.
+Numerical investigations are performed to demonstrate the performance of the proposed pGP approach.
+In particular, we compare the prediction accuracy between the proposed method and the existing ones, illustrate the uncertainty quantification capabilities and how the pGP can help to identify important parameters.
+Experiment setup and data.
+Consider, $\mathcal{A}({\bm{\theta}})x(t,\bm{s})=0$, an advection-diffusion process, where $x(t,\bm{s})$ is a spatio-temporal process in space $\bm{s}\in\mathbb{S}\subset\mathbb{R}^{2}$ and time $t\in[0,T]$, and $\mathcal{A}({\bm{\theta}})$ is the advection-diffusion operator parameterized by $\bm{\theta}$:
+$\mathcal{A}x:=\dot{x}+\bm{\vec{v}}^{T}\triangledown x-\triangledown\cdot[\bm{D}\triangledown x]$ (32)
+where $\bm{\vec{v}}$, $\bm{D}$, $\triangledown$ and $\triangledown\cdot$ represent the velocity, diffusivity, decay, gradient and divergence, respectively.
+In this experiment, $\mathbb{S}$ is a 2D rectangular domain.
+To investigate the performance of adapting the POD basis using the proposed pGP, we first let $\bm{v}=(1,0)^{T}$ and $\bm{D}=\mathrm{diag}\{d_{1},d_{2}\}$, and generate data by letting $(d_{1},d_{2})$ take values from a 2D mesh grid with uniformly spaced $x$- and $y$-coordinates in the interval $[0.01,0.05]$ with a step size of 0.002 (more parameters will be included in subsequent investigations).
+Hence, a number of $21\times 21=441$ parameter settings are considered, and the process is simulated for each parameter setting.
+For illustrative purposes, Figure 5 shows two simulated processes respectively for two parameter settings $[0.01,0.01]$ and $[0.05,0.05]$ of $(d_{1},d_{2})$, and Figure 6 shows the four leading POD modes of the two processes.
+It is seen that the POD modes vary as the parameters $d_{1}$ and $d_{2}$ change, suggesting the potential need to adapt the POD basis for parameters change.
+Figure 5: Two simulated processes respectively at parameter settings $[0.01,0.01]$ (top) and $[0.05,0.05]$ (bottom) of $(d_{1},d_{2})$ .
+Figure 6: The four leading POD modes obtained from the simulated processes at parameter settings $[0.01,0.01]$ (top) and $[0.05,0.05]$ (bottom) of $(d_{1},d_{2})$.
+Performance metrics. To comprehensively evaluate the accuracy of the predicted POD bases, four performance metrics are adopted in our investigations:
+[…]
+[…]
+$\bullet$ Following the $\ell_{2}$ error above, we also report the relative error so as to assess the percentage increase of the $\ell_{2}$ error from the optimal to the predicted POD basis,
+$e_{R}(\bm{\theta})=\frac{e_{F}(\bm{\theta})-e^{*}(\bm{\theta})}{e^{*}(\bm{\theta})}\times 100\%$ (34)
+[…]
+$\bullet$ In addition, motivated by the subspace angle interpolation of POD bases, the third performance metric is obtained by computing the geometric distance, i.e., the principal angel, between the subspaces respectively spanned by the optimal POD basis and the predicted POD basis.
+$e_{A}(\bm{\theta})=\left(\sum_{i=1}^{r}\text{arccos}^{2}(\sigma_{i})\right)^{\frac{1}{2}}$ (35)
+[…]
+$\bullet$ Finally, we evaluate the worst case scenario by computing the $\ell_{\infty}$ prediction error as follows:
+[…]
+where $||\cdot||_{\infty}$ is the matrix infinity norm which shows the largest error among the columns of the matrix $\bm{D}(\bm{\theta})-\hat{\bm{\Phi}}(\bm{\theta})\hat{\bm{\Phi}}(\bm{\theta})^{T}\bm{D}(\bm{\theta})$.
+Numerical results.
+The data are divided into a training set and a testing set.
+The training set includes 36 parameter settings $\bm{\theta}=(d_{1},d_{2})\in\Theta^{\text{train}}$ from a coarse 2D mesh grid with uniformly spaced $x$- and $y$-coordinates in the interval $[0.01,0.05]$ with a larger step size of 0.008.
+Data generated from the remaining 405 parameter settings, $\bm{\theta}\in\Theta^{\text{test}}$, are reserved for testing.
+Figure 7 shows the four types errors, $e_{F}$, $e_{R}$, $e_{A}$ and $e_{I}$, for all 405 testing cases while letting the reduced dimension to be $r=5$.
+Here, three different approaches are used, including the POD interpolation (using a Gaussian radial kernel), the global POD approach, and the proposed pGP approach.
+It is noted that, the proposed pGP approach yields the best performance in all four performance metrics.
+In particular,
+$\bullet$ In terms of $e_{F}$ and $e_{R}$, the proposed pGP outperforms the interpolation approach and the global POD approach respectively in 365 and 395 out of the 405 testing cases.
+$\bullet$ In terms of $e_{A}$, the proposed pGP outperforms the interpolation approach and the global POD approach in 351 and 385 out of the 405 testing cases, respectively.
+$\bullet$ In terms of $e_{I}$, the proposed pGP outperforms the interpolation approach and the global POD approach in 372 and 349 out of the 405 testing cases, respectively.
+Figure 7: Testing errors, $e_{F}$, $e_{R}$, $e_{A}$ and $e_{I}$, for all 405 testing cases ($r=5$)
+Figure 8: Testing errors, $e_{F}$, $e_{R}$, $e_{A}$ and $e_{I}$, for all 405 testing cases ($r=10$)
+Next, we increase the reduced dimension to $r=10$, and the updated comparison results are shown in Figure 8.
+The proposed pGP still yields the best overall performance.
+We also note that, when the dimension of the reduced-order basis increases, all three approaches yield smaller errors in terms of $e_{F}$, $e_{A}$ and $e_{I}$.
+However, in terms of the relative error $e_{R}$, the errors seem to become larger.
+This is not surprising as further investigation shows that the first 10 optimal POD modes (obtained from performing SVD directly on the testing snapshot data) can already capture about 97% of the total energy, making the denominator $e^{*}$ in (34) very small.
+Figure 9: The proposed pGP includes the global POD approach as its special case by yielding almost identical $e_{F}$ errors for all 405 testing cases when the dimension of the reduced-order basis is increased to $r=20$.
+Relationship between the proposed and global POD approaches.
+Following the discussions above, one may wonder what if the dimension of the reduced-order basis is further increased.
+Note that, an important property of the proposed pGP is that it includes the use of global POD basis as its special case.
+As already discussed in Sec.3.4, consider the squared Exponential kernel, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\{-(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})/2\}$, where $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{4}^{-2}\}$.
+If the estimated $\xi_{1}$, $\xi_{3}$ and $\xi_{4}$ are all very small, then, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ is close to zero and it is implied by (28) that $\bm{u}_{*}=\bm{0}$ and $)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}$)).
+Because the global POD basis is chosen as the reference point, any subspace (corresponding to different $\bm{\theta}\in\Theta^{\text{test}}$) on the Grassmann manifold can be essentially seen as the i.i.d.
+samples from a pGD, $)\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}$)).
+Hence, the point prediction, $\hat{p}_{*}$, is close to the subspace spanned by the global POD basis.
+This theoretical result is illustrated in Figure 9, which shows that the proposed pGP yields almost identical $e_{F}$ errors as the global POD approach does for all 405 testing cases when the dimension of the reduced-order basis is increased to $r=20$.
+In the Appendix, we provide one more example that shows how the proposed pGP includes the use of global POD as its special case.
+Figure 10: Bootstrapping standard deviation of the predicted subspace over the parameter space: the uncertainty becomes larger for parameter settings away from training points indicated by “++”.
+[…]
+To illustrate the output generated by the procedure above, Figure 10 shows the bootstrapping standard deviation of the predicted subspace at each testing parameter setting over the parameter space (here, $\tilde{n}=1000$).
+In this figure, the small cross “++” indicates the training points where there is no uncertainty.
+A clear pattern is observed: the uncertainty is the lower for parameter settings closer to the training points, while the uncertainty becomes larger for parameter settings away from training points.
+In general, the uncertainty is the largest halfway between existing training points.
+Such a pattern is consistent with our intuition.
+Parameter Selection Capabilities.
+As discussed in Sec.3.4, it is often the case that not all parameters are equally important in terms of their predictive power for POD basis modes.
+In other words, the POD modes can be sensitive to some parameters while independent of the others.
+Note that, this idea may seem to share some similarities with the concept of active subspace (Constantine et al.,, 2014).
+The difference between the concept of active subspace and parameter selection discussed here is that, the former identifies a low-dimensional space of model input which contains the detected directions of the strongest variability in model output, while the latter refers to the process of identifying the parameters which have the strongest predictive power of the POD modes in a statistical model (for a specific problem and given a specific training dataset).
+To demonstrate the parameter selection capabilities, we expand the parameter space and consider the change of both the velocity $\bm{v}=(v_{1},v_{2})$ and diffusivity $\bm{D}=\text{diag}\{d_{1},d_{2}\}$.
+In particular, the training data are simulated for $(v_{1},v_{2})$ values from a 2D mesh grid with uniformly spaced x- and y-coordinates in the interval [-0.1, 0.1] with a step size of 0.05, and for $(d_{1},d_{2})$ values from a 2D mesh grid with uniformly spaced x- and y-coordinates in the interval [0.01, 0.05] with a step size of 0.05.
+Before the pGP model is trained, the values of each of the four parameters are standardized between 0 an 1, and the standardized parameter space is $0,1]^{4}$.
+In Rasmussen and Williams, (2005), the squared Exponential kernel is considered when introducing the parameter selection problem for GP.
+In our experiment, we adopt the same type of kernel as follows with a minor modification:
+$\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}.$ (37)
+where $\bm{\theta}=(v_{1},v_{2},d_{1},d_{2})$, $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{3}^{-2},\xi_{4}^{-2},\xi_{4}^{-2}\}$, and $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ are the parameters in the kernel function to be estimated from data.
+Note that, $\xi_{3}$ and $\xi_{4}$ are known as the characteristic length-scales for velocity and diffusivity parameters.
+Following the idea of automatic relevance determination, a smaller length-scale indicates a stronger effect of a parameter on the prediction (Neal,, 2012).
+One may also note that we let $v_{1}$ and $v_{2}$ share the same length-scale $\xi_{3}$, and let $d_{1}$ and $d_{2}$ share the same length-scale $\xi_{4}$.
+This set up is appropriate considering the parameter mesh grids used to generate the training data, and there is no strong reason to justify why $v_{1}$ and $v_{2}$ (or, $d_{1}$ and $d_{2}$) should have very different length-scales.
+Figure 11: Growth of the estimated length-scales $\hat{\xi}_{3}$ and $\hat{\xi}_{4}$ as the tuning parameter $\gamma$ increases.
+The parameters $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ are estimated by the maximum likelihood method (Rasmussen and Williams,, 2005).
+In particular, to demonstrate the parameter selection capabilities, an $\ell_{2}$ regularization is imposed on $\xi_{3}$ and $\xi_{4}$, i.e., $\gamma(\xi_{3}^{-2}+\xi_{4}^{-2})$ for some tuning parameter $\gamma\geq 0$.
+Note that, (i) the regularization is only imposed on the length-scales $\xi_{3}$ and $\xi_{4}$, and (ii) as $\gamma$ increases, the regularization tends to make both $\xi_{3}$ and $\xi_{4}$ larger, i.e., to make the parameters (features) less influential in the regression model.
+Hence, the length-scale associated with the parameter that has less predictive power is expected to grow faster, helping us identify the important parameters.
+Figure 11 shows the estimated length-scales $\hat{\xi}_{3}$ and $\hat{\xi}_{4}$ for five different values of $\gamma$, i.e., 0, 500, 1000, 1500 and 2000.
+It is seen that, (i) when $\gamma=0$ (i.e., no regularization is added), the estimated values of $\xi_{3}$ and $\xi_{4}$ are respectively $\hat{\xi}_{3}=0.249$ (length-scale associated with the velocity parameter) and $\hat{\xi}_{4}=0.208$ (length-scale associated with the diffusion parameter).
+Because $\hat{\xi}_{3}$ is slightly larger than $\hat{\xi}_{4}$, indicating that the diffusion parameter may have a slightly higher predictive power than the velocity parameter; (ii) when $\gamma$ increases, both $\hat{\xi}_{3}$ and $\hat{\xi}_{4}$ become larger as expected; (iii) when $\gamma=2000$, the estimated length scale $\hat{\xi}_{3}=1.199$ is about 2.6 times higher than $\hat{\xi}_{4}=0.456$, clearly indicating that the diffusion parameter has a higher predictive power than the velocity parameter over the parameter space considered in this experiment.
+The observations above demonstrate the parameter selection capabilities of the proposed pGP approach when constructing a predictive model for the POD modes.
+5 Conclusions
+This work successfully demonstrated the great potential of introducing machine learning to the field of robust reduced-order modeling.
+A supervised learning problem, based on the proposed pGP regression, has been developed to predict the optimal POD basis.
+The proposed approach leverages an injective mapping from the parameter space to the Grassmann manifold that contains the optimal vector subspaces.
+The condition for such a mapping to be injective has been obtained, and sufficient technical details were presented for predicting the POD basis using the proposed pGP regression.
+Numerical experiments showed that the proposed pGP approach not only improves the accuracy of ROM against parameter change, but also enables uncertainty quantification and parameter selection, potentially reducing the dimension of the parametric space when adapting the POD basis.
+We would also like to point out that there exists a critical link between the proposed pGP and statistical experimental design/reinforcement learning that enable sequential exploration of a system’s behavior in a reduced parameter space.
+This link is expected to create important future research trajectories, potentially attracting more researchers into this field and making tangible contributions to Scientific Machine Learning research and its applications.
+Code and data are available on GitHub: https://github.com/dnncode/pGP.
+Appendix A Heat process in powder bed fusion
+In the Appendix, we provide one more example that shows how the proposed pGP includes the use of global POD basis as its special case, and yields more accurate predictions of the POD basis than interpolation.
 Experiment setup.
-In Example I, we consider the flow passing over a cylinder in a rectangle domain.
-The Lattice Boltzmann method is used to simulate the flow, i.e., to generate the training data for our model [22].
-In particular, the Bhatnagar–Gross–Krook (BGK) approximation to the Navier-Stokes equation is:
-(32) $\dot{f}+(\bm{v}\cdot\triangledown)f=-\tau(f-\bar{f})$
-where $f(\bm{x})$ be the phase-space density of fluids at location $\bm{x}$, $\bm{v}$ is the velocity, $\tau$ is the timescale of which collisions happen, and $\bar{f}$ represents the equilibrium state.
-In (32), the left side $(\frac{\partial}{\partial t}f+(\bm{v}\cdot\triangledown))f$ represents the streaming of particles, while the right side $-\tau(f-\bar{f})$ represents the collision.
-The model assumes that the fluid locally relaxes to equilibrium over a characteristic timescale $\tau$ related to the kinematic viscosity.
-The larger the $\tau$ is, the larger is the kinematic viscosity.
-Figure 5: Simulated flow velocity at times ($\tau=0.53$)
-The fluid equation (32) is solved by the Lattice Boltzmann method using the Python code in [22].
-As shown in Figure 5, we consider a $400\times 100$ rectangle domain.
-A solid cylinder is placed at location $(100,50)$ with a radius of 13.
-The velocity goes from left to right with a magnitude of $v=2.3$ per time step, and 5000 time steps are simulated.
-The initial condition follows a Gaussian distribution with mean zero and standard deviation 0.01.
-Reflective boundary conditions are considered for the cylinder, the top and the bottom boundaries, while the fluid is absorbed by the right boundary.
-Figure 5 shows the velocity of the flow, with $\tau=0.53$, at times 1, 2000, 4000 and 5000.
-The Von Karman vortices are visible at times 4000 and 5000.
-Figures 6 and 7 show the 1st, 4th, 7th and 10th POD modes for $\tau=0.53$ and $\tau=0.7$, respectively.
-We see that the POD modes change as the parameter $\tau$ changes, suggesting the necessity of adapting POD basis.
-Figure 6: The 1st, 4th, 7th and 10th POD modes for $\tau=0.53$
-Figure 7: The 1st, 4th, 7th and 10th POD modes for $\tau=0.7$
-To evaluate the accuracy of the predicted ROM through cross validation, we generate data for 33 values of $\tau$ ranging from 0.53 to 0.7 with a step size of 0.005.
-Then, the data are divided into a training and a testing sets.
-The training set includes the data generated from the following 11 parameter settings,
-$\tau^{\text{train}}=\{0.53,0.545,0.56,0.575,0.59,0.605,0.62,0.635,0.65,0.665,0.68\},$
-while the data generated from the remaining 22 parameter settings are used for testing.
-In particular, we let $m(\tau)$ yields a constant basepoint which is the subspace spanned by the global POD basis.
-The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$.
-Note that, $\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}$ does not depend on $\sigma_{\bm{K}}$ when $\bm{K}=\sigma_{\bm{K}}\bm{I}$, and only $\xi_{1}$ and $\xi_{2}$ are to be estimated by minimizing the cross-validation error.
-We obtain $\hat{\xi}_{1}\approx 1$ and $\hat{\xi}_{2}\approx 0.5$.
-Experiment Results.
-Figure 8 shows the out-of-sample testing errors for the proposed approach, the existing POD interpolation method, and the use of global POD bases.
-The errors are measured by the $\ell_{2}$ Frobenious norm defined under (2).
-Some important observations are summarized as follows:
-$\bullet$ The proposed approach yields much smaller $\ell_{2}$ error than the existing POD interpolation method from testing cases #1 to #6, as well as from #19 to #22 (i.e., towards the two boundaries of the parameter space of $\tau$).
-This consist observation suggests that the existing Lagrangian interpolation may have some stability issues for parameters near the boundaries of the parameter space, as investigated in [12].
-In addition, as discussed in the introduction section, the proposed approach allows us to optimally estimate the model parameters based on data, while the coefficients of the Lagrangian interpolation are independent from data.
-Overall, the proposed approach yields smaller $\ell_{2}$ error than the existing POD interpolation method for 21 out of the 22 testing cases (except for testing case #14 where the $\ell_{2}$ errors are almost identical).
-$\bullet$ The benefit of interpolating POD basis over the use of global POD bases is also demonstrated.
-The proposed approach yields smaller $\ell_{2}$ errors than the use of the global POD bases for 17 out of the 22 testing cases.
-Figure 8: Comparison of out-of-sample testing errors measured by the Frobenious norm
-Figure 9: Standard deviation associated with the ROM at time 100.
-As a statistical learning approach, the proposed approach is capable of quantifying the uncertainty of the output generated from the ROM.
-Figure 9 shows the bootstrapping standard deviation of the output of the ROM for testing cases #1, #8, #15 and #22.
-Within each iteration, a subspace is randomly sampled from the distribution $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}))$ obtained from (28), and an ROM is constructed on that sampled space.
-This procedure is repeated for 1000 times, and the standard deviation of the ROM output at time 100 is computed for each spatial location.
-4.2 Experiment II: Powder Bed Fusion Additive Manufacturing
-Powder Bed Fusion (PBF) with metals or polymers is one of the most common 3D printing technologies in industrial additive manufacturing (AM).
+We focus on the transient heat transfer in the single-track L-PBF (laser Powder Bed Fusion) process as shown in Figure 12.
+PBF with metals or polymers is one of the most common 3D printing technologies in industrial additive manufacturing (AM).
 PBF works by jointing powdered material point by point using energy sources such as lasers or electron beams.
-Thermal analysis is essential for product quality control which is key to spreading laser-based AM techniques [33] and manufacturing process design [17].
-Figure 10: Illustration of problem setup, spatial domain and boundary conditions.
-Experiment Setup.
-In this example, we focus on the transient heat transfer in the single-track L-PBF (laser PBF) process as shown in Figure 10.
-The vacuum chamber is preheated to a specific temperature and the spatial domain is a combination of the solid substrate and the powder layer.
-The transient heat transfer process is governed by a PDE:
-(33) $\rho c_{p}\frac{\partial u(x,t)}{\partial t}-\nabla\cdot k\nabla u(x,t)=0,\quad(x,t)\in\Omega\times[0,t_{f}],$
+Thermal analysis is essential for product quality control which is key to spreading laser-based AM techniques (Yan et al.,, 2018) and manufacturing process design (Liao et al.,, 2022).
+The vacuum chamber is preheated to a specific temperature and the spatial domain is a combination of the solid substrate and the powder layer. The transient heat transfer process is governed by a PDE:
+$\rho c_{p}\frac{\partial u(x,t)}{\partial t}-\nabla\cdot k\nabla u(x,t)=0,\quad(x,t)\in\Omega\times[0,t_{f}],$ (38)
 where $\rho$ denotes the density, $c_{p}$ is the specific heat capacity, $k$ is the thermal conductivity of the material, $u(x,t)$ represents the desired temperature over a space-time domain $\Omega\times[0,t_{f}]$.
 The heat flux boundary condition in PBF manufacturing processes is described as
-(34) $(-k\nabla u(x,t))\cdot\mathbf{n}=q_{h}+q_{c}+q_{r},\quad(x,t)\in\partial\Omega\times[0,t_{f}],$
+$(-k\nabla u(x,t))\cdot\mathbf{n}=q_{h}+q_{c}+q_{r},\quad(x,t)\in\partial\Omega\times[0,t_{f}],$ (39)
 where $\mathbf{n}$ is the (outward) unit normal vector to the surface, $q_{h}$, $q_{c}$, and $q_{r}$ are the heat flux due to the heat source, convection and radiation, respectively.
+Figure 12: Illustration of problem setup, spatial domain and boundary conditions.
 A moving Gaussian heat source considered as follows:
-(35) $q_{h}=-\frac{2\alpha\tau}{\pi\gamma^{2}}\text{exp}(-\frac{2\|x-x^{c}(t)\|^{2}}{\gamma^{2}}),\quad x\in\partial\Omega_{t},$
+$q_{h}=-\frac{2\alpha\tau}{\pi\gamma^{2}}\text{exp}(-\frac{2\|x-x^{c}(t)\|^{2}}{\gamma^{2}}),\quad x\in\partial\Omega_{t},$ (40)
 where $\partial\Omega_{t}$ denotes the top surface of domain, $\alpha$ is the heat absorptivity, $\tau$ is the laser power, $\gamma$ is the effective laser (or electron) beam radius, $x^{c}(t)$ is the center location of the energy beam at time $t$.
 The convection and radiation, for $x\in\partial\Omega_{t}\cup\partial\Omega_{s}$, are modeled as
-(36) $q_{c}(x,t)=h(u(x,t)-u_{a}),\quad q_{r}(x,t)=\sigma_{s}\epsilon(u(x,t)^{4}-u_{a}^{4}),$
+$q_{c}(x,t)=h(u(x,t)-u_{a}),\quad q_{r}(x,t)=\sigma_{s}\epsilon(u(x,t)^{4}-u_{a}^{4}),$ (41)
 where $\partial\Omega_{s}$ denotes the side surfaces of domain, $h$ is the heat convection coefficient, $u_{a}$ is the ambient temperature, $\sigma_{s}$ is the Stefan-Boltzmann constant, $\epsilon$ is the emissivity of the material.
@@
 Note that, $\partial\Omega=\partial\Omega_{t}\cup\partial\Omega_{s}\cup\partial\Omega_{b}$.
-The Finite Element method (FEM) is used to solve the governing equations. Multiplying both sides of (33) by a test function $\phi$ and integrating over $\Omega$ gives
-[…]
+The finite element method is used to solve the governing equations. Multiplying both sides of (38) by a test function $\phi$ and integrating over $\Omega$ gives
+$\int_{\Omega}\rho c_{p}\frac{\partial u(x,t)}{\partial t}\phi dV-\int_{\Omega}\nabla\cdot k\nabla u(x,t)\phi dV=0,\quad(x,t)\in\Omega\times[0,t_{f}],$ (42)
 The solution $u(x,t)$ and the test function are assumed to belong to Hilbert spaces.
@@
 According to Green’s first identity, (42) gives
-[…]
+$\int_{\Omega}\rho c_{p}\frac{\partial u(x,t)}{\partial t}\phi dV+\int_{\Omega}k\nabla u(x,t)\cdot\nabla\phi dV+\int_{\partial\Omega}(-k\nabla u(x,t))\cdot\mathbf{n}\phi dS=0,$ (43)
 Then, the discretization by Galerkin method is applied to the weak formulation and the approximated solution $u_{d}(x,t)$ is given by $u_{d}(x,t)=\sum_{i=1}^{N}u_{i}(t)\psi_{i}(x)$, where $N$ is the number of nodes.
 For every test function $\psi_{j}(x)$, we have
-[…]
+$\rho c_{p}\sum_{i=1}^{N}\frac{\partial u_{i}(t)}{\partial t}\int_{\Omega}\psi_{i}\psi_{j}dV+k\sum_{i=1}^{N}u_{i}(t)\int_{\Omega}\nabla\psi_{i}\nabla\psi_{j}dV+\sum_{i=1}^{N}\int_{\partial\Omega}(-k\nabla u_{i}(t)\psi_{i})\cdot\mathbf{n}\psi_{j}dS=0$ (44)
 Figure 13: Temperature field at 600W laser power with the heat source moving at a speed of 5mm/s.
 We generate data for 7 laser power levels (in Watt): $\{500,550,600,650,700,750,800\}$, using the finite element method.
-As an illustration, Fig.11 shows the temperature field at selected times.
-Since seven laser power levels are used to generate the data, a Leave-One-Out Cross-Validation (LOOCV) is performed to compare the model performance.
-Similar to Example 1, we let $r=10$ and $m(\tau)$ be a constant basepoint which is the subspace spanned by the global POD basis.
+As an illustration, Figure 13 shows the temperature field at selected times.
+Since seven laser power levels are used to generate the data, the Leave-One-Out Cross-Validation (LOOCV) is performed to compare the model performance.
+We let $r=10$ and $m(\tau)$ be a constant basepoint which is the subspace spanned by the global POD basis.
 The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$.
@@
 $\bullet$ More importantly, the proposed pGP is able to provide an explanation to such an observation.
-In short, this is because the proposed pGP (i) involves the use of global POD basis as its special case (due to its parameter selection capabilities), and (ii) successfully identifies that the ROM, in this example, is not sensitive to the change of parameters, and the use of the global POD basis becomes appropriate.
+In short, this is because the proposed pGP (i) involves the use of global POD basis as its special case (due to its parameter selection capabilities), and (ii) successfully identifies that the POD basis, in this example, is not sensitive to the change of parameters, and the use of the global POD basis becomes appropriate.
+Figure 14: Leave-One-Out Cross-Validation (LOOCV) errors.
 To elaborate, the kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$ in this example.
-When LOOCV is performed, we note that the estimated values for $\xi_{2}$ ranges between 40 to 80, effectively making $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ close to zero.
-In this case, (28) immediately implies that $\bm{u}_{*}=\bm{0}$ and $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}))$.
+When the LOOCV is performed, we note that the estimated values for $\xi_{2}$ ranges between 40 to 80, effectively making $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ close to zero.
+In this case, (28) immediately implies that $\bm{u}_{*}=\bm{0}$ and $)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}$)).
 Because the global POD basis is chosen as the reference point, then, any subspace (corresponding different power levels $\tau$) on the Grassmann manifold can be seen as i.i.d.
-samples from a PGD, $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}))$.
+samples from a pGD, $)\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}$)).
 Hence, the point prediction, $\hat{p}_{*}$, is close to the subspace spanned by the global POD basis, justifying the observation in Figure 14.
-Figure 12: Leave-One-Out Cross-Validation (LOOCV) errors.
-4.3 Experiment III: Advection-Diffusion Processes
-Advection-diffusion equations are widely used to describe the transfer of particles, energy, or other physical quantities inside a physical system due to diffusion and advection.
-Experiment Setup.
-We consider a general form of advection-diffusion process $\mathcal{A}({\bm{\theta}})x(t,\bm{s})=0$, where $x(t,\bm{s})$ is a spatio-temporal process in space $\bm{s}\in\mathbb{S}\subset\mathbb{R}^{2}$ and time $t\in[0,T]$, and $\mathcal{A}({\bm{\theta}})$ is the advection-diffusion operator parameterized by $\bm{\theta}$:
-(40) $\mathcal{A}x:=\dot{x}+\bm{\vec{v}}^{T}\triangledown x-\triangledown\cdot[\bm{D}\triangledown x]$
-[…]
-Let $\bm{x}(t;\bm{\theta})=(x(t,\bm{s}_{1};\bm{\theta}),\cdots,x(t,\bm{s}_{n};\bm{\theta}))$ be an $n$-dimensional vector representing the discretized solutions at a set of locations $\{\bm{s}_{1},\bm{s}_{2},\cdots,\bm{s}_{n}\}\in\mathbb{S}$, we represent the advection-diffusion process by a dynamical system that is linear in state
-(41) $\bm{x}(t;\bm{\theta})=\bm{F}\bm{\alpha}(t;\bm{\theta}),\quad\dot{\bm{\alpha}}(t;\bm{\theta})=\bm{G}(\bm{\theta})\bm{\alpha}(t;\bm{\theta})$
-Here, $\bm{F}$ is an $n\times n$ matrix with its $i$-th row being $\bm{f}(\bm{s}_{i})=(f_{1}(\bm{s}_{i}),...,f_{n}(\bm{s}_{i}))$ where $f_{j}(\bm{s})=\exp(i\bm{k}_{j}^{T}\bm{s})$ is the Fourier basis function and $\bm{k}_{j}$ is the spatial wavenumber, $\bm{\alpha}(t)\in\mathbb{R}^{n}$ is the state vector, $\bm{G}(\bm{\theta})$ is an $n\times n$ state transition matrix operator, and the initial condition is denoted by $\bm{x}_{0}$.
-Note that, the matrix $\bm{G}(\bm{\theta})$ depends on the parameters in (40) and is given by [19].
-In projection-based ROM of parametric dynamic systems[6], we seek a ROM of the form to replace the original large-scale model (41) :
-(42) $\bm{x}_{r}(t;\bm{\theta})=\bm{F}_{r}\bm{\alpha}_{r}(t;\bm{\theta}),\quad\dot{\bm{\alpha}}_{r}(t;\bm{\theta})=\bm{G}_{r}(\bm{\theta})\bm{\alpha}_{r}(t;\bm{\theta})$
-where $\bm{\alpha}_{r}(t)$ is the $r$-dimensional reduced-order state vector ($r\ll n$), $\bm{x}_{r}(t)$ is an approximation of $\bm{\alpha}_{r}(t)$, and $\bm{F}_{r}\in\mathbb{R}^{n\times r}$ and $\bm{G}_{r}\in\mathbb{R}^{r\times r}$.
-Following the Galerkin’s method, it is known that $\bm{\alpha}_{r}(t)=\bm{\Phi}_{\bm{\alpha}}^{T}\bm{\alpha}(t)$, $\bm{G}_{r}=\bm{\Phi}_{\bm{\alpha}}^{T}\bm{G}\bm{\Phi}_{\bm{\alpha}}$ and $\bm{F}_{r}=\bm{F}\bm{\Phi}_{\bm{\alpha}}$, where $\bm{\Phi}_{\bm{\alpha}}$ is a matrix basis with orthogonal columns.
-Given any parameter $\bm{\theta}$, we obtain from the snapshot data matrix $\bm{D}(\bm{\theta})=[\bm{x}(t_{1};\bm{\theta});\bm{x}(t_{2};\bm{\theta});\cdots;\bm{x}(t_{n_{T}};\bm{\theta})]$ the POD basis $\bm{\Phi}$.
-Because $\bm{x}(t;\bm{\theta})=\bm{F}\bm{\alpha}(t;\bm{\theta})$, i.e., $\bm{x}(t;\bm{\theta})$ is the left unitary transformation of $\bm{\alpha}(t;\bm{\theta})$, the invariance property of SVD to unitary transformation yields $\bm{\Phi}_{\bm{\alpha}}=\bm{F}{\bm{\Phi}}$[16].
-To evaluate the accuracy of adapting ROM using the proposed pGP, we let $\bm{v}=(v_{x},0.005)^{T}$, $\bm{D}=\mathrm{diag}\{0.001,0.001\}$ and generate data by letting $v_{x}$ vary from -0.02 to 0.02 with a step size of 0.001.
-The same initial condition is used for all $v_{x}$ values.
-In total, a number of 42 snapshot data matrices are generated for 42 values of $v_{x}$.
-The data are divided into a training and a testing sets.
-The training set includes the data generated from the following 11 parameter settings, $v_{x}^{\text{train}}=\{-0.016,-0.012,-0.008,-0.004,-0.002,0,0.004,0.008,0.012,0.016,0.02\}$, while the data generated from the remaining 31 parameter settings are used for testing.
-In the testing phase, the trained pGP model is used to predict the optimal POD basis $\bm{\Phi}_{\bm{\alpha}}$ for a given value of $v_{x}$, and the corresponding ROM (42) is constructed.
-Then, we solve the ROM to obtain the reduced-order solution $\bm{x}_{r}(t)$ and re-construct the full-order solutions by $\bm{x}(t)=\bm{\Phi}\bm{x}_{r}(t)$.
-Experiment Results.
-To show and compare the performance of different approaches, Figure 13 shows the (re-constructed) solution from the proposed pGP, (re-constructed) solution from the existing POD interpolation approach, (re-constructed) solution using the global POD, and the (ground truth) solution by directly solving the full-order system.
-For illustrative purposes, these figures show the results for $v_{x}=0.006$ (which is not included in the training dataset) at times 1 (the initial condition), 5, 10, and 15.
-We see that the accuracy of the (reconstructed) solutions obtained from the existing POD interpolation approach deteriorates starting from time 5, while the proposed pGP approach provides more accurate solutions.
-(a) t=1 (initial condition)
-(b) t=5
-(c) t=10
-(d) t=15
-Figure 13: Comparison between the proposed pGP (row 1 of each subfigure), the existing POD interpolation approach (row 2), the use of global POD (row 3), and the (ground truth) solution by directly solving the full-order system (row 4) at $v_{x}=0.006$
-To further show the advantages of the proposed pGP approach, Figure 14 presents the comparison of the overall accuracy for all testing cases. The improvement by the proposed pGP is clearly observed.
-(a) $\ell_{2}$ error
-(b) $L_{\infty}$ error
-Figure 14: Comparison between the proposed approach, existing POD interpolation approach and the use of global POD for constructing ROM against parameter changes.
-4.4 Code and Data
-Code and data are submitted as Supplementary Materials.
-Because the size of the dataset is about 3.5GB, users may download the data from the following link: https://drive.google.com/file/d/1Ub_TnaB0--cW3tcNlSsr2qiVdliH2xdJ/view?usp=sharing.
-5 Conclusions
-This work successfully demonstrated the great potential of introducing machine learning to the field of robust reduced-order modeling.
-A supervised learning problem, based on the proposed pGP regression, has been developed to predict the optimal POD basis.
-The proposed approach leverages an injective mapping from the parameter space to the Grassmann Manifold that contains the optimal vector subspaces.
-The condition for such a mapping to be injective has been obtained, and sufficient technical details were presented for predicting POD basis using the proposed pGP regression.
-Numerical experiments showed that the proposed pGP approach not only improves the accuracy of ROMs against parameter change, but also enables uncertainty quantification and parameter selection, potentially reducing the dimension of the parametric space when adapting ROMs.
-We would also like to point out that there exists a critical link between the proposed pGP and statistical experimental design/reinforcement learning that enable sequential exploration of a system’s behavior in a reduced parameter space.
-This link is expected to create important future research trajectories, potentially attracting more researchers into this field and making tangible contributions to Scientific Machine Learning research and its applications.